### v1

# Words as Trigger Points in Social Media Discussions

Abstract
Trigger points are a concept introduced by Mau, Lux, and Westheuser (2023) to study qualitative focus group interviews and understand polarisation in Germany. When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals react affectively to such triggers and show strong and negative emotional responses. In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analysing a large amount of social media posts. We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics. We find that such trigger words affect user engagement and have noticeable consequences on animosity in online discussions. We share empirical evidence of trigger words causing animosity, and how they provide incentives for hate speech, adversarial debates, and disagreements. Our work is the first to introduce trigger points to computational studies of online communication. Our findings are relevant to researchers interested in online harms and who examine how citizens debate politics and society in light of affective polarisation.
1 Introduction
Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations. This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to negative reactions, hate speech, and abusive language in online debates.
Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
Trigger points are related to how individuals perceive themselves in society (Mau, Lux, and Westheuser 2023). People have subjective expectations about acceptable behaviour such as explicit knowledge of law-like rules and, more importantly, implicit understanding of appropriate actions. These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable in individuals understanding of the social contract. Trigger points jeopardise what individuals understand as the fabric of society and their own role in it.
Mau, Lux, and Westheuser (2023) propose four different mechanisms why trigger points affect individuals.
  • Inequality. When expectations about equality might not be met. Either similar people are treated unequally, or, to the contrary, people who should be treated differently are treated the same. For example, social benefits might be at eye level with minimal wage earners.
  • Norm transgression. It is the transgression of what is considered “normal” behaviour. A typical example relates to the lavish lifestyle of the privileged.
  • Slippery slope. There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it. “Opening the borders so that a country could be flooded with immigrants” is an example in this context.
  • Behavioural demands. When society may be perceived as demanding unreasonable behaviour. The role of pronouns is a point in case where the generic masculine pronouns “he/him” might trigger some, and the neutral pronoun “they/them” might trigger others.
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality (Mau, Lux, and Westheuser 2023). In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
  • Hypothesis 1 more user engagement, which incites a higher number of messages;
  • Hypothesis 2 higher levels of animosity, which may lead to polarisation, negativity, anger, and hate.
In our study we focus on how individual words can already act as trigger points in online debates. We look into five trigger words (Rwanda, Brexit, NHS, Vaccine, Feminism) in the context of UK politics and collect >>>100 million related Reddit posts. We then leverage textual and Natural Language Processing (NLP) tools to look into the extent to which these words lead to potentially negative, emotional, and harmful responses, and analyse the results based on a systematic comparison between three treatment and control groups:
  1. 1.
    Space: specific subreddits we expect to be triggered by the selected words.
  2. 2.
    Semantics: words similar in meaning to the target words but without the triggering component.
  3. 3.
    Time: time periods identified as relevant for each specific trigger word.
Our results show that these trigger words cause higher levels of engagement and more animosity, measured as an increase in the controversiality, negative sentiment, anger, and hate speech in a thread.
2 Related Work
Trigger points (Mau, Lux, and Westheuser 2023) bear great potential to understand polarisation and division in online discussions. However, the existence or validity of the concept has not been tested on a large scale. Trigger points can be particularly relevant to work on social media. They help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and reasons for their escalation (De Kock and Vlachos 2022; van der Meer et al. 2023). Further to triggering ad hominem attacks, escalation can lead to hate speech. There is, however, no unanimous agreement on how it differs from abusive, and offensive language (Basile et al. 2019; Davidson et al. 2017; Talat and Hovy 2016).
Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al. 2019; Sap et al. 2019; ElSherief et al. 2021; Ocampo et al. 2023), and dogwhistles, which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper 2021), we know little about the sources of these phenomena. Examining what can trigger harmful reactions may help us make the detection process more robust and avoid the well-known problem of selection bias (Wiegand, Ruppenhofer, and Kleinbauer 2019; Ousidhoum, Song, and Yeung 2020) caused by the common data collection strategies based on hashtags (Talat and Hovy 2016) and keywords (Davidson et al. 2017).
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al. 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Consequently, studying trigger points can help us track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al. 2021).
3 Methodology
Mau, Lux, and Westheuser (2023) developed the concept of trigger points based on qualitative interviews with focus groups to analyse the perceived polarisation of German society. We use this concept to study online communication on social media at a large scale. To critically assess its reliability, we test whether already individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
3.1 Term selection
We selected terms likely to provoke reactions from individuals in the United Kingdom. Specifically, we include National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism in our study. These trigger words were chosen because they are associated with high-profile public debates, policy changes, and social movements, which have been known to provoke significant public engagement and emotional responses. Table 1 indicates the respective mechanism through which the terms may trigger individuals.
Rwanda Brexit NHS Vaccine Feminism
Inequality
Norm transgression
Slippery Slope
Behavioural Demands
Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
Rwanda
Rwanda plays a particular role in the UK’s immigration policy. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country. However, with the then government, the term’s meaning changed in early 2022 and became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. Instead of handling asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be processed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent (The Economist 2023). With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly relating to concerns about “unfair favours towards migrants and unregulated immigration”.
Brexit
The UK’s departure from the European Union has been controversial since its inception. Even eight years post-exit, debates persist regarding the soundness of the decision. The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega 2018). Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving 2019). Additionally, misinformation and disinformation related to Brexit have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
NHS
Due to recent financial austerity measures, the UK government was accused of neglecting the health care system and reducing its budget, which leads to deteriorating services (The Guardian 2022; The Independent 2024). As a result, the NHS has become a hot topic of discussion among the UK population. Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
Vaccine
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing. Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi 2021). Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021). Moreover, the proliferation of misinformation and conspiracy theories regarding vaccines has further fuelled public division (Hayawi et al. 2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
Threads Comments Year
Rwanda 0.05 1.87 2020, 2022
Brexit 0.37 11.57 2020 - 2022
NHS 0.34 6.20 2020 - 2022
Vaccine 1.42 28.11 2019, 2022
Feminism 0.94 17.79 2020 - 2022
EU 0.2 12.4 2020 - 2022
Red Cross 0.02 0.2 2020 - 2022
SubSaharan 0.08 1.5 2020, 2022
Surgery 0.5 3.9 2019, 2022
Woman 2.1 19.3 2020 - 2022
Table 2: Numbers, in millions, of threads, comments, and time period for each trigger word.
Feminism.
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019). Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al. 2022).  [^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al. 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
Keyword Subreddits
Feminism
AskMen, MensRights,
antifeminists, AndrewTateTalk
Brexit
neoliberal, brexit, ireland, unitedkingdom,
The_Donald, ukpolitics, tories,
NHS
neoliberal, CoronavirusUK, unitedkingdom,
ukpolitics, tories, labourUK
Rwanda
neoliberal, tories, labourUK, currenteventsuk
ukpolitics, uknews, unitedkingdom
Vaccine
conspiracy, neoliberal, unpopularopinion,
antifascistsofreddit, coronavirusuncensored,
CoronavirusUK, Coronavirus, wuhan_flu
Table 3: Subreddits selected for each trigger word.
Feminist Rwanda Brexit NHS Vaccine
Be Af DiD CI Be Af DiD CI Be Af DiD CI Be Af DiD CI Be Af DiD CI
Treatment
Only
47 53 45 55 46 54 46 54 47 53
Space 49 51 4.1 (3.8, 4.5) 50 50 9.7 (8.2, 11.3) 49 51 6.3 (6.1, 6.6) 49 51 6.1 (5.7, 6.4) 49 51 4.1 (3.5, 3.9)
Semantic 48 52 4.6 (3.9, 5.2) 50 50 9.8 (6.6, 14.9) 47 53 3.2 (2.6, 3.7) 53 47 13.4 (8.7, 18.1) 49 51 5.0 (3.1, 6.9)
Temporal - 49 51 7.2 (2.9, 11.6) - - 50 50 5.9 (4.1, 7.6)
Table 4: Percentage of comments per thread before and after each trigger appearance (Hypothesis 1). Difference in Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 7.
3.2 Data collection
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al. 2020). We searched for exact matches of specific keywords in comments, ignoring the respective case, and gathered all corresponding posts from 2019 to 2022. Subsequently, the entire discussion thread for each comment extracted was collected [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.. We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al. 2017). Posts marked as “[deleted]” by the API were also removed. Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point. As a final filtering step, and considering the fast-paced nature of Reddit, we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word. Table 2 shows the total number of comments, threads, and comments with triggers collected for each trigger point and the time period they were gathered.
3.3 Research design
We identify the causal effect of trigger points with the difference-in-difference estimator (Angrist and Pischke 2009; Card and Krueger 1994). This estimator is based on systematically comparing a treatment and a control group. It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words.
Control groups
We systematically compare treatment and control groups along three different dimensions.
1. Space
Discussions on Reddit are structured in different fora, also called subreddits. Aiming to investigate the different effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 3). We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
2. Semantics
We also compare trigger words with a term that shares a similar meaning without going against the deeper underlying transgressions that incite trigger responses. To members of the subreddits in Table 3, using an explicit trigger word or its more harmless control term should make a real difference. We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^4] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia. to Rwanda.
3. Time
In the third experiment, we make use of the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time. For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022. For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic). Again, we expect that users from the subreddits in Table 3 are triggered by those only in 2022 but not in the respective earlier years.
Statistical analysis
We use the difference-in-difference (DiD) estimator to analyse the causal effect of the trigger word on different online behaviours. The idea behind the DiD estimator is to compare the outcome change between the treatment and control groups before and after the trigger word. Specifically, we calculate the difference in the average outcome change for the treatment group before and after the trigger word and then subtract the difference in the average outcome change for the control group. If the treatment has a causal effect, we expect a larger change for the treatment group than for the control group. For each thread, we normalise by the actual number of messages, and calculate the share of messages before and after the appearance of the trigger word. To estimate the causal effect, we specify the DiD as an OLS regression (Angrist and Pischke 2009):
[…]
where proportion is the per-thread-normalised target feature and the interaction term is given by
$\text{interaction}=\text{trigger\_dummy}\times\text{treatment\_dummy}.interaction = trigger_dummy × treatment_dummy$ .
3.4 Features
We annotate several features to examine our initial hypotheses and study the impact of trigger words in online discussions. We simply use post counts for the first hypothesis (H1) related to user engagement. The second hypothesis (H2) requires a measure of animosity. To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al. 2019) and are trained on social media corpora, specifically X (Twitter). Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, unstructured text. Hence, we expect the trained models to perform well in either context (Priya et al. 2019).
User engagement (Hypothesis 1)
Our first Hypothesis suggests that trigger words lead to higher user engagement. We simply rely on the number of posts to investigate this. That is, we expect to see an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
Animosity (Hypothesis 2)
Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, the presence a negative sentiment, anger, and different forms of hate speech.
Controversiality
We examine whether a comment is controversial using the Reddit controversiality feature. Reddit offers a measure that identifies a comment as controversial if it receives similar amounts of upvotes and downvotes. This metric is based on the users’ approval, so care must be taken when considering already controversial subreddits. We expect trigger words to cause more controversiality.
Figure 2: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
Sentiment analysis
For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al. 2022) which identifies a post as negative, neutral, or positive. This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019). The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al. 2021), politics (Guimaraes et al. 2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022). We expect trigger words to cause more negative sentiments.
Negative emotions
We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al. 2022) to assign one or more emotions to each tweet. This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al. 2018), covering 11 different emotions [^5] Detailed list of emotions can be found in Appendix A.1). For our analysis we focus on anger specifically, and we expect to observe more anger after a trigger word.
Figure 3: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
Hate speech
Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados 2023) to detect hate speech. It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^6] Detailed list of categories can be found in Appendix A.2). The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al. 2021), women’s rights (Farrell et al. 2019), and health issues (Walter et al. 2022). Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
4 Results
In this section, we present the results for the two hypotheses tested in this paper.
4.1 User engagement (Hypothesis 1)
Our initial hypothesis is that if trigger words significantly influence online discussions, there should be more messages after mentioning these terms. The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 4).
In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all the terms we selected. It is largest with Rwanda where we observe a 10%-points (Table 4) increase. To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group. In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate. The lowest value is a 3.2%-point increase for the semantic comparison in Brexit. The largest increase is for NHS in the semantic control group, where we observe a 13.4% point increase.
When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent. For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space related experiment. It shows the overall number of messages across all threads. This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial. Similar patterns are also observed in the semantic and temporal control groups for Rwanda (Appendix B, Figure 9).
Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our initial hypothesis on the effect of trigger words on online discussions.
Trigger Setting Feature DiD CI
Brexit Space Contr 1.0 (0.8, 1.3)
N.Sent 3.3 (3.0, 3.5)
Anger 2.7 (2.4, 3.0)
Hate -0.1 (-0.5, 0.4)
Semantics Contr 1.4 (0.7, 2.1)
N.Sent 2.6 (2.0, 3.1)
Anger 3.8 (2.1, 5.5)
Hate -0.5 (-1.9, 0.9)
Feminist Space Contr 0.3 (-0.2, 0.8)
N.Sent 2.6 (2.2, 3.0)
Anger 1.4 (1.0, 1.9)
Sexism 0.6 (0.1, 1.1)
Semantics Contr 0.1 (-0.8, 1.1)
N.Sent 3.0 (2.3, 3.6)
Anger 1.4 (1.0, 1.9)
Sexism 1.4 (0.5, 2.3)
NHS Space Contr 1.1 (0.7, 1.4)
N.Sent 3.1 (2.8, 3.5)
Anger 2.7 (2.3, 3.1)
Hate -0.3 (-0.9, 0.3)
Semantics Contr 4.5 (-3.2, 12.2)
N.Sent 3.5 (-1.2, 8.3)
Anger 4.2 (-11.6, 20.0)
Hate -1.4 (-12.5, 9.6)
Trigger Setting Feature DiD CI
Rwanda Space Contr 2.3 (0.8, 3.7)
N.Sent 5.5 (4.0, 6.9)
Anger 4.0 (2.7, 5.3)
Racism 1.8 (0.0, 3.6)
Semantics Contr 6.6 (0.2, 13.0)
N.Sent 5.5 (1.5, 9.6)
Anger 5 (-4.6, 14.6)
Racism 14.6 (-0.7, 29.9)
Temporal Contr 5.9 (0.5, 11.3)
N.Sent 4.8 (0.8, 8.8)
Anger 2.8 (-1.0, 6.6)
Racism 0.9 (-11.1, 12.8)
Vaccine Space Contr 1.5 (1.3, 1.8)
N.Sent 3.0 (2.8, 3.2)
Anger 2.0 (1.7, 2.2)
Hate 0.4 (0.0, 0.8)
Semantics Contr 2.2 (1.1, 3.4)
N.Sent 5.4 (4.7, 6.1)
Anger 4.4 (1.6, 7.2)
Hate -0.7 (-2.4, 0.9)
Temporal Contr -0.3 (-3.0, 2.3)
N.Sent 2.9 (1.4, 4.4)
Anger 5.5 (1.4, 9.6)
Hate -2.0 (-5.3, 1.3)
Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
4.2 Animosity (Hypothesis 2)
Our second hypothesis suggests that trigger words may increase the frequency of polarising messages and induce animosity in the discussion.
Figure 4: Average controversiality per thread for for the treatment and control (space) groups of Rwanda.
Controversiality.
We analyse the average controversiality of each thread and compare the proportion of controversial threads before and after the introduction of the trigger words. Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word. Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups with the biggest difference observed on the semantic comparison for Rwanda, with a causal effect of 6.6% points.
Figure 5: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
However, significant differences in controversiality do not appear across all settings. Notably, there seems to be no increase in controversial comments when considering Feminism, and in the temporal and semantic settings for Vaccine and NHS, respectively.
Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda in both the space control and treatment groups, before and after its occurrence. The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points. A similar trend is seen in both the semantic and temporal control settings, as documented in Appendix B, Figure 8).
(a) Rwanda
(b) Feminism
Figure 6: Distribution of hateful comments in the treatment and control groups (Space) of the trigger points Rwanda and Feminism.
Sentiment & emotion.
We further investigate the effect trigger words have on sentiment and emotion. Based on our hypothesis, trigger words cause users to express more negatively charged emotions. In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
When considering the negatively charged comments of the treatment and control groups for each trigger word, Table 5 indicates a clear trend. We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS. The biggest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word within both the space and semantics control groups.
Let us consider the vaccine trigger, Figure 2, as an example. Here, we again plot the overall number of messages we observe. Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned. However, the treatment group shows more messages than the control group, suggesting that introducing the trigger words affects the expression of negative sentiment.
Similarly to the increase in negative sentiment, trigger words systematically cause more anger. We observe a positive causal effect in all conditions, with only 3 non-significant effects. Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison). As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 5. This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
Hate speech.
As a final indicator of animosity, we also consider hate speech. In our analysis, we use any kind of hate speech for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism. As seen from Table 5, the impact of trigger words regarding the occurrence of hateful comments is not as prevalent as anger and negative sentiment. We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
We exemplify our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in. We observe a distinct pattern, particularly in the context of Rwanda and Feminism where the volume of hateful content increases post-trigger. Figure 6 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control counterparts in regards to space.
5 Conclusion
In this paper, we introduce the concept of trigger words, and study how they affect social media conversations. We reported empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine). In short, we find that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. We consistently observe more posts for all trigger words and across all control settings. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examined four different features that can indicate such effects. Controversiality, negative sentiment, and anger all show strong evidence that trigger words influence the discussion along those lines. We also investigated the causal effect of trigger words on hate speech. Where we could rely on a model that is capable of identifying domain-specific forms of hate speech—Sexism in the case of Feminism and Racism in the case of Rwanda—we find evidence that suggest trigger words play a causal role, here, too.
6 Limitations
Despite potentially more accurate results, we refrained from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al. 2023).
Selecting words as trigger points and the subreddits for analysing them offers room for improvement. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community.
The goal of this paper is to verify trigger words as a concept. As such, we identify and manually select a few words as candidates according to our own expertise and social science literature. If we do not find evidence for the effects of trigger words in these easy cases, it would not make sense to study trigger words and their effects any further.
7 Ethics Statement
All the data utilised in our research is sourced from publicly available information or collected using the official Reddit API. All the information is provided in an aggregated fashion, without reporting sensitive information from individual users.
The automatic methods used can be abused by those with the power to repress people with opposing opinions. Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use. Researchers interested in replicating our experiments can contact us directly.
Appendix A Model Categories
A.1 Emotion Categories
The twitter-roberta-base-emotion-multilabel model classifies each entry in one or more of the following classes: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust.
A.2 Hate Speech Categories.
The twitter-roberta-base-hate-multiclass model classifies each entry in one of the following classes: not_hate, sexism, racism, religion, other, sexual_orientation, disability
Appendix B Extended Results
Figure 7 displays the number of comments posted before and after the introduction of each trigger point on the data collected.
Figure 8 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
Figure 9 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
Figure 10 displays the distribution of negative comments, based on their sentiment, posted before and after the introduction of the Brexit, Feminism, Rwanda, and NHS trigger points in the Space control setting.
(a) Brexit
(b) Feminism
(c) NHS
(d) Rwanda
(e) Vaccine
Figure 7: Absolute count of posts before and after each trigger point.
(a) Temporal Setting
(b) Semantic Setting
Figure 8: Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
(a) Semantic Comparison
(b) Temporal Comparison
Figure 9: Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
(a) Brexit
(b) Feminism
(c) Rwanda
(d) NHS
Figure 10: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
Appendix C Computational Resources
An NVIDIA GeForce RTX 4090 GPU was utilised for the experiments conducted. 170 hours spent for the inference process of sentiment analysis, emotion detection, and hate speech detection models.

### v2

@@
-Words as Trigger Points in Social Media Discussions
-Abstract
-Trigger points are a concept introduced by Mau, Lux, and Westheuser (2023) to study qualitative focus group interviews and understand polarisation in Germany.
+Words as Trigger Points in Social Media Discussions—A Large-Scale Case Study about UK Politics on Reddit
+Abstract.
+Trigger points are a concept introduced by Mau et al.
+(2023) to study qualitative focus group interviews and understand polarisation in Germany.
 When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned.
-In the original studies, individuals react affectively to such triggers and show strong and negative emotional responses.
-In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analysing a large amount of social media posts.
+In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned.
+In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analyzing a large amount of social media posts.
 We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics.
-We find that such trigger words affect user engagement and have noticeable consequences on animosity in online discussions.
-We share empirical evidence of trigger words causing animosity, and how they provide incentives for hate speech, adversarial debates, and disagreements.
+We find that such trigger words affect user engagement and have noticeable consequences how online discussions occur.
+Trigger words cause animosity and incentivise hate speech, adversarial debates, and disagreements.
 Our work is the first to introduce trigger points to computational studies of online communication.
-Our findings are relevant to researchers interested in online harms and who examine how citizens debate politics and society in light of affective polarisation.
-1 Introduction
-Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions.
+The findings are relevant to researchers interested in affective computing, online deliberation, and those examining how citizens debate politics and society in light of affective polarisation.
+Affective Computing, Polarization, Reddit, Trigger Words, User Engagement
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENCE 2025; 28 April–2 May, 2025; Sydney, Australiaccs: Information systems World Wide Webccs: Computing methodologies Natural language processing
+1. Introduction
+Trigger points, introduced by Mau et al.
+(2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions.
 Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations.
-This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to negative reactions, hate speech, and abusive language in online debates.
-Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
-Trigger points are related to how individuals perceive themselves in society (Mau, Lux, and Westheuser 2023).
-People have subjective expectations about acceptable behaviour such as explicit knowledge of law-like rules and, more importantly, implicit understanding of appropriate actions.
+This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
+Figure 1. Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
+Trigger points are related to how individuals perceive themselves in society.
+People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions.
 These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world.
 People feel triggered if these normative understandings are questioned.
-More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable in individuals understanding of the social contract.
-Trigger points jeopardise what individuals understand as the fabric of society and their own role in it.
-Mau, Lux, and Westheuser (2023) propose four different mechanisms why trigger points affect individuals.
+More formally, Mau et al.
+(2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract.
+Trigger points jeopardise what individuals consider as the fabric of society and their role in it.” Trigger points can affect individuals through four mechanisms.
   • Inequality.
@@
 Going against deeply held beliefs about individuals and their societal role causes discomfort.
-People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality (Mau, Lux, and Westheuser 2023).
-In the social media context, we expect that this has two direct consequences (Figure 1).
-Trigger points lead to:
-  • Hypothesis 1 more user engagement, which incites a higher number of messages;
-  • Hypothesis 2 higher levels of animosity, which may lead to polarisation, negativity, anger, and hate.
-In our study we focus on how individual words can already act as trigger points in online debates.
-We look into five trigger words (Rwanda, Brexit, NHS, Vaccine, Feminism) in the context of UK politics and collect >>>100 million related Reddit posts.
-We then leverage textual and Natural Language Processing (NLP) tools to look into the extent to which these words lead to potentially negative, emotional, and harmful responses, and analyse the results based on a systematic comparison between three treatment and control groups:
-  1. 1.
-    Space: specific subreddits we expect to be triggered by the selected words.
-  2. 2.
+People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality.
+In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
+  • Hypothesis 1 more user engagement, which incites a higher number of messages or comments;
+  • Hypothesis 2 higher levels of animosity, which incite polarisation, negativity, anger, and hate.
+Our study focuses on how individual words can act as trigger points in online debates.
+We look into five words (Rwanda, Brexit, NHS, Vaccine, Feminism) we expect to work as trigger points in the context of UK politics.
+Collecting >>>100 million related Reddit posts, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses.
+The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
+  1. (1)
+    Space: specific subreddits where the selected words trigger users.
+  2. (2)
     Semantics: words similar in meaning to the target words but without the triggering component.
-  3. 3.
+  3. (3)
     Time: time periods identified as relevant for each specific trigger word.
-Our results show that these trigger words cause higher levels of engagement and more animosity, measured as an increase in the controversiality, negative sentiment, anger, and hate speech in a thread.
-2 Related Work
-Trigger points (Mau, Lux, and Westheuser 2023) bear great potential to understand polarisation and division in online discussions.
+Our results show that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
+2. Related Work
+Trigger points (Mau et al., 2023) bear great potential to understand polarisation and division in online discussions.
 However, the existence or validity of the concept has not been tested on a large scale.
 Trigger points can be particularly relevant to work on social media.
-They help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and reasons for their escalation (De Kock and Vlachos 2022; van der Meer et al.
-2023).
+They help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and reasons for their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023).
 Further to triggering ad hominem attacks, escalation can lead to hate speech.
-There is, however, no unanimous agreement on how it differs from abusive, and offensive language (Basile et al.
-2019; Davidson et al.
-2017; Talat and Hovy 2016).
-Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al.
-2019; Sap et al.
-2019; ElSherief et al.
-2021; Ocampo et al.
-2023), and dogwhistles, which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper 2021), we know little about the sources of these phenomena.
-Examining what can trigger harmful reactions may help us make the detection process more robust and avoid the well-known problem of selection bias (Wiegand, Ruppenhofer, and Kleinbauer 2019; Ousidhoum, Song, and Yeung 2020) caused by the common data collection strategies based on hashtags (Talat and Hovy 2016) and keywords (Davidson et al.
-2017).
-Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al.
-2019), Blodgett et al.
+There is, however, no unanimous agreement on how it differs from abusive and offensive language (Basile et al., 2019; Davidson et al., 2017; Talat and Hovy, 2016).
+Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al., 2019; Sap et al., 2019; ElSherief et al., 2021; Ocampo et al., 2023), and dog-whistles which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper, 2021), we know little about the sources of these phenomena.
+Examining what can trigger harmful reactions may help us make more informed choices when collecting data, thus, leading to more robust detection models and mitigating the well-known problem of selection bias (Wiegand et al., 2019; Ousidhoum et al., 2020).
+Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al.
 (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model.
-Consequently, studying trigger points can help us track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al.
-2021).
-3 Methodology
-Mau, Lux, and Westheuser (2023) developed the concept of trigger points based on qualitative interviews with focus groups to analyse the perceived polarisation of German society.
-We use this concept to study online communication on social media at a large scale.
-To critically assess its reliability, we test whether already individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
-3.1 Term selection
+Studying trigger points helps track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al., 2021).
+Further, understanding how trigger words affect online communication also sheds light on how language affects polarisation on social media (Karjus and Cuskley, 2024).
+It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
+3. Methodology
+We use the UK context as a large-scale case study to observe polarisation when individuals communicate on social media.
+Our goal is to report behavioural mechanisms similar to those observed by Mau et al.
+(2023) when conducting interviews to analyse the perceived polarisation of the German society.
+Based on a curated set of likely candidates, we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
+3.1. Term selection
 We selected terms likely to provoke reactions from individuals in the United Kingdom.
@@
 Behavioural Demands
-Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
-Rwanda
+Table 1. Trigger words and their respective trigger mechanisms (described in the introduction).
+Rwanda.
 Rwanda plays a particular role in the UK’s immigration policy.
 [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country.
-However, with the then government, the term’s meaning changed in early 2022 and became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership.
-Instead of handling asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be processed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent (The Economist 2023).
-With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly relating to concerns about “unfair favours towards migrants and unregulated immigration”.
-Brexit
+However, in early 2022 the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership.
+Instead of processing asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be assessed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent.
+With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
+Brexit.
 The UK’s departure from the European Union has been controversial since its inception.
 Even eight years post-exit, debates persist regarding the soundness of the decision.
-The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega 2018).
-Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving 2019).
-Additionally, misinformation and disinformation related to Brexit have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022).
+The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega, 2018).
+Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving, 2019).
+Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022).
 Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
-NHS
-Due to recent financial austerity measures, the UK government was accused of neglecting the health care system and reducing its budget, which leads to deteriorating services (The Guardian 2022; The Independent 2024).
+NHS.
+Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the health care system (The Guardian, 2022; The Independent, 2024).
 As a result, the NHS has become a hot topic of discussion among the UK population.
-Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015).
+Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick et al., 2015).
 This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
-Vaccine
+Vaccine.
 Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years.
 Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing.
-Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi 2021).
-Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021).
-Moreover, the proliferation of misinformation and conspiracy theories regarding vaccines has further fuelled public division (Hayawi et al.
-2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
+Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi, 2021).
+Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021).
+Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
 Threads Comments Year
@@
 Woman 2.1 19.3 2020 - 2022
-Table 2: Numbers, in millions, of threads, comments, and time period for each trigger word.
+Table 2. Numbers, in millions, of threads, comments, and time period for each trigger word.
 Feminism.
-Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019).
-Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al.
-2022).
+Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz, 2019).
+Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al., 2022).
 [^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women.
-This has led to the formation of online echo chambers (Farrell et al.
-2019) where such views are amplified and propagated.
+This has led to the formation of online echo chambers (Farrell et al., 2019) where such views are amplified and propagated.
 Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
@@
 CoronavirusUK, Coronavirus, wuhan_flu
-Table 3: Subreddits selected for each trigger word.
+Table 3. Subreddits selected for each trigger word.
 Feminist Rwanda Brexit NHS Vaccine
@@
 Temporal - 49 51 7.2 (2.9, 11.6) - - 50 50 5.9 (4.1, 7.6)
-Table 4: Percentage of comments per thread before and after each trigger appearance (Hypothesis 1).
-Difference in Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons.
+Table 4.
+Percentage of comments per thread before and after each trigger appearance (Hypothesis 1).
+Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons.
 Bold scores indicate statistically significant results.
 Detailed distribution of comments can be seen in Appendix B, see Figure 7.
-3.2 Data collection
-Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al.
-2020).
-We searched for exact matches of specific keywords in comments, ignoring the respective case, and gathered all corresponding posts from 2019 to 2022.
-Subsequently, the entire discussion thread for each comment extracted was collected [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume..
-We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al.
-2017).
+3.2. Data collection
+Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al., 2020).
+We searched for exact matches of specific keywords in comments from 2019 to 2022, and then collected the entire discussion thread for each extracted comment [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume..
+We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al., 2017).
 Posts marked as “[deleted]” by the API were also removed.
 Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
-As a final filtering step, and considering the fast-paced nature of Reddit, we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word.
-Table 2 shows the total number of comments, threads, and comments with triggers collected for each trigger point and the time period they were gathered.
-3.3 Research design
-We identify the causal effect of trigger points with the difference-in-difference estimator (Angrist and Pischke 2009; Card and Krueger 1994).
-This estimator is based on systematically comparing a treatment and a control group.
+As a final filtering step we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word.
+Previous research has shown that Reddit posts typically receive the majority of interactions within a short window after the post is created (Thukral et al., 2018).
+Moreover, early responses are often important in shaping the direction of the conversation (Singer et al., 2016; Horne et al., 2017).
+Considering the fast-paced nature of Reddit, the 30 minutes window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data.
+Table 2 shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered.
+3.3. Research design
+We identify the causal effect of trigger points with the difference-in-difference estimator (DiD) which systematically compares a treatment with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994).
 It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group.
 In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words.
-Control groups
+There should be a more substantive behavioural change for the treatment group than for the control group.
+3.3.1. Control groups
 We systematically compare treatment and control groups along three different dimensions.
+[^4] Identifying the causal effect in three completely unrelated comparisons helps address concerns about eventual endogeneity.
 1. Space
@@
 2. Semantics
-We also compare trigger words with a term that shares a similar meaning without going against the deeper underlying transgressions that incite trigger responses.
+We also compare trigger words with a term that shares a similar meaning, but that does not go against the deeper underlying transgressions responsible for trigger responses.
 To members of the subreddits in Table 3, using an explicit trigger word or its more harmless control term should make a real difference.
@@
 3. Time
-In the third experiment, we make use of the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time.
+In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time.
 For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022.
@@
 Again, we expect that users from the subreddits in Table 3 are triggered by those only in 2022 but not in the respective earlier years.
-Statistical analysis
-We use the difference-in-difference (DiD) estimator to analyse the causal effect of the trigger word on different online behaviours.
-The idea behind the DiD estimator is to compare the outcome change between the treatment and control groups before and after the trigger word.
-Specifically, we calculate the difference in the average outcome change for the treatment group before and after the trigger word and then subtract the difference in the average outcome change for the control group.
-If the treatment has a causal effect, we expect a larger change for the treatment group than for the control group.
-For each thread, we normalise by the actual number of messages, and calculate the share of messages before and after the appearance of the trigger word.
-To estimate the causal effect, we specify the DiD as an OLS regression (Angrist and Pischke 2009):
+3.3.2. Statistical analysis
+We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 4) and 48 experiments for H2 (Table 5).
+For each experiment, we prepare the data in the same way.
+The treatment/placebo is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread.
+To test H1, we count the number of posts for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group.
+We then calculate the share of posts before and after the intervention.
+For example, if there are 8 posts during the 30 minutes before the trigger word and 12 posts during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention.
+For experiments with H2, we count the number of the respective animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
+We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke, 2009):
 […]
@@
 $\text{interaction}=\text{trigger\_dummy}\times\text{treatment\_dummy}.interaction = trigger_dummy × treatment_dummy$ .
-3.4 Features
-We annotate several features to examine our initial hypotheses and study the impact of trigger words in online discussions.
+Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al., 2023).
+It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time.
+We investigate the plausibility of the assumption in section 4.3.
+3.4. Features
+We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions.
 We simply use post counts for the first hypothesis (H1) related to user engagement.
 The second hypothesis (H2) requires a measure of animosity.
-To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al.
-2019) and are trained on social media corpora, specifically X (Twitter).
-Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, unstructured text.
-Hence, we expect the trained models to perform well in either context (Priya et al.
-2019).
-User engagement (Hypothesis 1)
+To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al., 2019) and are trained on social media corpora, specifically X (Twitter).
+Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text.
+Hence, we expect the trained models to perform well in either context (Priya et al., 2019).
+3.4.1. User engagement (Hypothesis 1)
 Our first Hypothesis suggests that trigger words lead to higher user engagement.
@@
 That is, we expect to see an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
-Animosity (Hypothesis 2)
+3.4.2. Animosity (Hypothesis 2)
 Hypothesis 2 suggests that trigger words cause more animosity in a debate.
-We operationalise this with controversial comments, the presence a negative sentiment, anger, and different forms of hate speech.
+We operationalise this with controversial comments, the presence of negative sentiment, anger, and different forms of hate speech.
 Controversiality
@@
 We expect trigger words to cause more controversiality.
-Figure 2: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
+Figure 2. Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
 Sentiment analysis
-For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al.
-2022) which identifies a post as negative, neutral, or positive.
-This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019).
-The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al.
-2021), politics (Guimaraes et al.
-2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022).
+For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) which identifies a post as negative, neutral, or positive.
+This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019).
+The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender biases (Marjanovic et al., 2022).
 We expect trigger words to cause more negative sentiments.
 Negative emotions
-We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al.
-2022) to assign one or more emotions to each tweet.
-This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al.
-2018), covering 11 different emotions [^5] Detailed list of emotions can be found in Appendix A.1).
-For our analysis we focus on anger specifically, and we expect to observe more anger after a trigger word.
-Figure 3: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
+We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022) to assign one or more emotions to each tweet.
+This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018), covering 11 different emotions [^6] Detailed list of emotions can be found in Appendix A.1).
+For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
+Figure 3. Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
 Hate speech
-Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados 2023) to detect hate speech.
+Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados, 2023) to detect hate speech.
 It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^7] Detailed list of categories can be found in Appendix A.2).
-The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al.
-2021), women’s rights (Farrell et al.
-2019), and health issues (Walter et al.
-2022).
+The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al., 2021), women’s rights (Farrell et al., 2019), and health issues (Walter et al., 2022).
 Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
-4 Results
+4. Results
 In this section, we present the results for the two hypotheses tested in this paper.
-4.1 User engagement (Hypothesis 1)
+4.1. User engagement (Hypothesis 1)
 Our initial hypothesis is that if trigger words significantly influence online discussions, there should be more messages after mentioning these terms.
@@
 When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent.
-For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space related experiment.
+For instance, Figure 3 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment.
 It shows the overall number of messages across all threads.
@@
 Hate -2.0 (-5.3, 1.3)
-Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
-4.2 Animosity (Hypothesis 2)
+Table 5. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
+4.2. Animosity (Hypothesis 2)
 Our second hypothesis suggests that trigger words may increase the frequency of polarising messages and induce animosity in the discussion.
-Figure 4: Average controversiality per thread for for the treatment and control (space) groups of Rwanda.
+Figure 4. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
 Controversiality.
@@
 Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word.
-Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups with the biggest difference observed on the semantic comparison for Rwanda, with a causal effect of 6.6% points.
-Figure 5: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
+Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the semantic comparison for Rwanda—a causal effect of 6.6% points.
 However, significant differences in controversiality do not appear across all settings.
 Notably, there seems to be no increase in controversial comments when considering Feminism, and in the temporal and semantic settings for Vaccine and NHS, respectively.
-Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda in both the space control and treatment groups, before and after its occurrence.
+Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence.
 The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points.
@@
 (b) Feminism
-Figure 6: Distribution of hateful comments in the treatment and control groups (Space) of the trigger points Rwanda and Feminism.
+Figure 5. Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
 Sentiment & emotion.
@@
 We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS.
-The biggest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word within both the space and semantics control groups.
+The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and semantics experiments.
 Let us consider the vaccine trigger, Figure 2, as an example.
@@
 Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned.
-However, the treatment group shows more messages than the control group, suggesting that introducing the trigger words affects the expression of negative sentiment.
+However, the treatment group shows more messages than the control group, suggesting that the trigger words affects the expression of negative sentiment.
 Similarly to the increase in negative sentiment, trigger words systematically cause more anger.
@@
 Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison).
-As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 5.
+As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure LABEL:fig:emotion_space_vaccine.
 This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
 Hate speech.
-As a final indicator of animosity, we also consider hate speech.
-In our analysis, we use any kind of hate speech for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism.
-As seen from Table 5, the impact of trigger words regarding the occurrence of hateful comments is not as prevalent as anger and negative sentiment.
+As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of hate speech or abusive language for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism.
+As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment.
 We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
-We exemplify our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in.
-We observe a distinct pattern, particularly in the context of Rwanda and Feminism where the volume of hateful content increases post-trigger.
-Figure 6 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control counterparts in regards to space.
-5 Conclusion
-In this paper, we introduce the concept of trigger words, and study how they affect social media conversations.
-We reported empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine).
-In short, we find that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
+We illustrate our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in.
+We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger.
+Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
+4.3. Testing the Parallel Trends Assumption.
+The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time.
+While there are no formal guarantees for the assumption to hold, there are means to argue in its favour.
+The first argument is a theoretical one.
+In neither of our experimental designs, we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word.
+Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (semantics), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
+We implement event studies for each experiment to further empirically support our argument.
+The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention.
+Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel.
+If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al., 2023).
+Figure 6. Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
+For each of our overall 60 experiments, we divide the 30 minutes before and after the first mention of a trigger word into six distinct time intervals $\{[-30,-20[,[-20,-10[,[-10,0[,[0,10[,[10,20[,[20,30[\}${ [ - 30 , - 20 [ , [ - 20 , - 10 [ , [ - 10 , 0 [ , [ 0 , 10 [ , [ 10 , 20 [ , [ 20 , 30 [ } instead of just two.
+This allows us to follow, in particular, the pre-treatment trends at a higher granularity.
+We calculate each discussion thread’s share of (hostile) messages in the six intervals.
+Averaging over all these data points per period, we subtract the control group from the treatment group and calculate the difference of each observation from the first observation period.
+Figure 6 charts the time-wise means over all 60 event studies.
+If the parallel trends assumption holds, the lines before the treatment at time = 0 should all remain on the horizontal line where the treatment effect = 0.
+Indeed, this is largely what we find.
+The mean treatment effect is close to 0 before the trigger word is mentioned for the first time.
+After the intervention, it is clearly above the horizontal 0-line.
+[^8] See Appendix for figures that include individual event studies for all experiments from H1 and H2.
+5. Conclusion
+In this paper, we present the concept of trigger words and study how they affect social media conversations in UK politics.
+We report empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine).
+In short, our results show that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
 Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word.
-We consistently observe more posts for all trigger words and across all control settings.
-Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examined four different features that can indicate such effects.
-Controversiality, negative sentiment, and anger all show strong evidence that trigger words influence the discussion along those lines.
-We also investigated the causal effect of trigger words on hate speech.
-Where we could rely on a model that is capable of identifying domain-specific forms of hate speech—Sexism in the case of Feminism and Racism in the case of Rwanda—we find evidence that suggest trigger words play a causal role, here, too.
-6 Limitations
-Despite potentially more accurate results, we refrained from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries).
+We consistently observe more posts across all trigger words and control settings.
+Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects.
+Trigger words cause more controversiality, more negative sentiment, and more anger.
+We also investigate the causal effect of trigger words on hate speech.
+Where we can, we rely on a model that is capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—.
+Our findings similarly confirm that trigger words play a causal role in this aspect as well.
+6. Limitations
+We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries).
 Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments.
-Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al.
-2023).
-Selecting words as trigger points and the subreddits for analysing them offers room for improvement.
+Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al., 2023).
+Selecting words as trigger points and their corresponding subreddits offers room for improvement.
 The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community.
-The goal of this paper is to verify trigger words as a concept.
-As such, we identify and manually select a few words as candidates according to our own expertise and social science literature.
-If we do not find evidence for the effects of trigger words in these easy cases, it would not make sense to study trigger words and their effects any further.
-7 Ethics Statement
-All the data utilised in our research is sourced from publicly available information or collected using the official Reddit API.
-All the information is provided in an aggregated fashion, without reporting sensitive information from individual users.
-The automatic methods used can be abused by those with the power to repress people with opposing opinions.
-Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use.
-Researchers interested in replicating our experiments can contact us directly.
+Similarly, we only investigate a very select set of potential trigger words in the context of the UK.
+However, our study had different goal: We consider our work to be a pilot study that intends to investigate whether the trigger points exist in online settings at all.
+Based on the empirical evidence we find, we can be confident to have found relevant empirical evidence that warrants further investigation.
+7. Ethics Statement
+All data utilised in our research is sourced from publicly available information or collected using the official Reddit API.
+All information is provided in an aggregated fashion without reporting sensitive information from individual users.
 Appendix A Model Categories
-A.1 Emotion Categories
+A.1. Emotion Categories
 The twitter-roberta-base-emotion-multilabel model classifies each entry in one or more of the following classes: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust.
-A.2 Hate Speech Categories.
+A.2. Hate Speech Categories.
 The twitter-roberta-base-hate-multiclass model classifies each entry in one of the following classes: not_hate, sexism, racism, religion, other, sexual_orientation, disability
@@
 (e) Vaccine
-Figure 7: Absolute count of posts before and after each trigger point.
+Figure 7. Absolute count of posts before and after each trigger point.
 (a) Temporal Setting
 (b) Semantic Setting
-Figure 8: Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
+Figure 8. Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
 (a) Semantic Comparison
 (b) Temporal Comparison
-Figure 9: Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
+Figure 9. Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
 (a) Brexit
@@
 (d) NHS
-Figure 10: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
-Appendix C Computational Resources
+Figure 10. Distribution of negatively charged comments in the Treatment and Control (Space) groups.
+Appendix C Extended Test of Parallel Trends Assumption
+(a) Experiments from H1.
+(b) Experiments from H2.
+Figure 11. Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
+Appendix D Computational Resources
 An NVIDIA GeForce RTX 4090 GPU was utilised for the experiments conducted. 170 hours spent for the inference process of sentiment analysis, emotion detection, and hate speech detection models.

### v3

@@
-Words as Trigger Points in Social Media Discussions—A Large-Scale Case Study about UK Politics on Reddit
-Abstract.
-Trigger points are a concept introduced by Mau et al.
-(2023) to study qualitative focus group interviews and understand polarisation in Germany.
-When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned.
+Words as Trigger Points in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit
+Abstract
+Political debates on social media sometimes flare up.
+From that moment on, users engage much more with one another; their communication is also more emotional and polarised.
+While it has been difficult to grasp such moments with computational methods, we suggest that trigger points are a useful concept to understand and ultimately model such behaviour.
+Established in qualitative focus group interviews to understand political polarisation (Mau, Lux, and Westheuser 2023), trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned.
 In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned.
-In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analyzing a large amount of social media posts.
-We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics.
-We find that such trigger words affect user engagement and have noticeable consequences how online discussions occur.
-Trigger words cause animosity and incentivise hate speech, adversarial debates, and disagreements.
-Our work is the first to introduce trigger points to computational studies of online communication.
-The findings are relevant to researchers interested in affective computing, online deliberation, and those examining how citizens debate politics and society in light of affective polarisation.
-Affective Computing, Polarization, Reddit, Trigger Words, User Engagement
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENCE 2025; 28 April–2 May, 2025; Sydney, Australiaccs: Information systems World Wide Webccs: Computing methodologies Natural language processing
-1. Introduction
-Trigger points, introduced by Mau et al.
-(2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions.
-Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations.
-This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
-Figure 1. Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
+Our paper finds that these trigger points also exist in online debates.
+We examine online deliberations on Reddit between 2020 and 2022 and collect $>$100 million comments from subreddits related to a set of words identified as trigger points in UK politics.
+Analysing the comments, we find that trigger words increase user engagement and animosity, i.e., more negativity, hate speech, and controversial comments.
+Introducing trigger points to computational studies of online communication, our findings are relevant to researchers interested in affective computing, online deliberation, and how citizens debate politics and society in light of affective polarisation.
+1 Introduction
+Online debates may suddenly escalate, and users intensely engage in emotional and confrontational deliberations.
+While these turning points are difficult to model and predict, they are essential moments in online communication.
+They mark the moment when a discussion is likely to get off the rails, and social media users stop debating with one another and instead debate against one another.
+Our study introduces trigger points as a useful concept to understand why and when social media users engage in such behaviour.
+Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions.
+They open the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
+Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
 Trigger points are related to how individuals perceive themselves in society.
 People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions.
-These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world.
+These deeply rooted beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world.
 People feel triggered if these normative understandings are questioned.
-More formally, Mau et al.
-(2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract.
-Trigger points jeopardise what individuals consider as the fabric of society and their role in it.” Trigger points can affect individuals through four mechanisms.
+More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract.
+Trigger points jeopardise what individuals consider the fabric of society and their own role in it.” The authors suggest that trigger points can affect individuals through four mechanisms.
   • Inequality.
@@
 There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it.
-“Opening the borders so that a country could be flooded with immigrants” is an example in this context.
+Claims such as “Opening the borders so that a country could be flooded with immigrants” is an example.
   • Behavioural demands.
@@
 Going against deeply held beliefs about individuals and their societal role causes discomfort.
-People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality.
+People respond to such triggers in an affective behavioural mode, increasing a debate’s intensity and emotionality.
 In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
-  • Hypothesis 1 more user engagement, which incites a higher number of messages or comments;
-  • Hypothesis 2 higher levels of animosity, which incite polarisation, negativity, anger, and hate.
-Our study focuses on how individual words can act as trigger points in online debates.
-We look into five words (Rwanda, Brexit, NHS, Vaccine, Feminism) we expect to work as trigger points in the context of UK politics.
-Collecting >>>100 million related Reddit posts, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses.
+  • Hypothesis 1: more user engagement, which leads to more messages or comments.
+  • Hypothesis 2: higher levels of animosity, which causes polarisation, negativity, anger, and hate.
+Our study seeks to establish the concept of trigger points in online communication.
+The research design and case selection closely follow this logic and examine whether and how individual words can act as trigger points in online debates.
+To this end, we conduct a large-scale case study and select five words that are ex-ante likely to work as trigger points in the context of UK politics.
+Specifically, we look into the words Rwanda, Brexit, NHS, Vaccine, Feminism.
+Collecting $>$100 million related Reddit comments, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses.
 The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
-  1. (1)
+  1. 1.
     Space: specific subreddits where the selected words trigger users.
-  2. (2)
-    Semantics: words similar in meaning to the target words but without the triggering component.
-  3. (3)
-    Time: time periods identified as relevant for each specific trigger word.
-Our results show that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
-2. Related Work
-Trigger points (Mau et al., 2023) bear great potential to understand polarisation and division in online discussions.
-However, the existence or validity of the concept has not been tested on a large scale.
-Trigger points can be particularly relevant to work on social media.
-They help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and reasons for their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023).
-Further to triggering ad hominem attacks, escalation can lead to hate speech.
-There is, however, no unanimous agreement on how it differs from abusive and offensive language (Basile et al., 2019; Davidson et al., 2017; Talat and Hovy, 2016).
-Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al., 2019; Sap et al., 2019; ElSherief et al., 2021; Ocampo et al., 2023), and dog-whistles which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper, 2021), we know little about the sources of these phenomena.
-Examining what can trigger harmful reactions may help us make more informed choices when collecting data, thus, leading to more robust detection models and mitigating the well-known problem of selection bias (Wiegand et al., 2019; Ousidhoum et al., 2020).
-Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al.
+  2. 2.
+    Contextual: words that appear in a similar context to the target words but without the triggering component.
+  3. 3.
+    Time: periods identified as relevant for each specific trigger word.
+We find that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
+2 Related Work
+Existing work studies how particular political words or phrases lead to more engagement (Brady et al.
+2017; Hameleers et al.
+2021) and posts with more emotion and animosity (Friedman et al.
+2004; Martin et al.
+2021; Schroth, Bain-Chekal, and Caldwell 2005; Rho and Mazmanian 2020) in various online settings.
+However, they lack a joint conceptual and theoretical framework—something we aim to provide.
+Trigger points are fundamental to understanding polarisation and division in online discussions.
+Given that they are at the root of different conflictive behaviours later on, they help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and their escalation (De Kock and Vlachos 2022; van der Meer et al.
+2023) triggering explicit and implicit ad hominem attacks, hate speech, abusive or offensive language on social media (Talat and Hovy 2016; Davidson et al.
+2017; Basile et al.
+2019; Ousidhoum et al.
+2019; Sap et al.
+2019; ElSherief et al.
+2021; Ocampo et al.
+2023).
+Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al.
+2019), Blodgett et al.
 (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model.
-Studying trigger points helps track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al., 2021).
-Further, understanding how trigger words affect online communication also sheds light on how language affects polarisation on social media (Karjus and Cuskley, 2024).
-It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
-3. Methodology
-We use the UK context as a large-scale case study to observe polarisation when individuals communicate on social media.
-Our goal is to report behavioural mechanisms similar to those observed by Mau et al.
-(2023) when conducting interviews to analyse the perceived polarisation of the German society.
-Based on a curated set of likely candidates, we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
-3.1. Term selection
-We selected terms likely to provoke reactions from individuals in the United Kingdom.
-Specifically, we include National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism in our study.
-These trigger words were chosen because they are associated with high-profile public debates, policy changes, and social movements, which have been known to provoke significant public engagement and emotional responses.
-Table 1 indicates the respective mechanism through which the terms may trigger individuals.
+Studying trigger points helps track the origins of bias, assess whether a social media post can lead to toxic comments (Dahiya et al.
+2021), and how language affects polarisation on social media (Karjus and Cuskley 2024).
+It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon, Brady, and Van Bavel 2022).
+Related to trigger points are dog-whistles—coded expressions that carry two meanings: a seemingly harmless one, and another one with a hidden hateful or provocative connotation that only insiders can decipher (Breitholtz and Cooper 2021).
+In social sciences, the concept of dog whistle has its roots in the study of racial politics in the United States, leading politicians to develop more subtle ways of appealing to racial prejudices (López 2013) and has evolved to encompass a broader range of phenomena, ranging from perpetuated systemic racism (Filimon and and 2024), to more broadly communication strategies of political ambiguity in identity politics(Tolvanen 2024), such as religion (Perry 2023) or immigration (Filimon 2016).
+Trigger points are similar in their equivocality and potential to evoke emotional responses.
+Conversely, using a dog whistle is conscious, but there is typically no intent when relying on trigger points.
+Moreover, while dog whistles are rhetorical devices to hide the real intention from some and create a rally-around-the-flag effect, trigger points are not hidden communication and incite a strong sense of otherness towards those covered by the trigger.
+Research on the automation of identifying affective triggers in online discussions remains limited.
+For instance, Almerekhi et al.
+(2022) developed a framework to identify triggers for toxic behaviour within Reddit discussions.
+Similar work typically uses a restricted set of lexical features, primarily sentiment analysis, and often examines the level of engagement some topics may generate (Aldous, An, and Jansen 2023).
+Others, such as Beel et al.
+(2022), explicitly investigate highly controversial topics like abortion, climate change, and gun control, attempting to predict contentiousness within discussions, often heavily relying on Reddit comment upvote ratios as a proxy for contentiousness.
+Finally, some studies focus on online conversational tensions by predicting conversational derailment, considering the overall progression of a conversation rather than specific terms as causes (Zhang et al.
+2018; Hessel and Lee 2019).
+3 Methodology
+To empirically demonstrate that individual words can trigger online debates, we look into most likely cases in the context of UK politics.
+Based on a curated set of likely candidate words for the political context in the UK (see Section 3.1), we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
+Our goal is to report behavioural mechanisms similar to those observed by Mau, Lux, and Westheuser (2023) when conducting interviews to analyse the perceived polarisation.
+If we could not find these effects in these intuitive cases, we would falsify the theory for online contexts.
+3.1 Term selection
+We selected National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism.
+These terms are associated with high-profile public debates, policy changes, and social movements in the UK.
+Hence, they typically provoke significant public engagement and emotional responses.
+We expect them to cause the intended reactions in British social media users.
+Table 1 indicates mechanisms through which the terms may trigger individuals.
 Rwanda Brexit NHS Vaccine Feminism
@@
 Behavioural Demands
-Table 1. Trigger words and their respective trigger mechanisms (described in the introduction).
+Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
 Rwanda.
-Rwanda plays a particular role in the UK’s immigration policy.
-[^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country.
-However, in early 2022 the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership.
-Instead of processing asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be assessed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent.
-With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
+Until late 2021, Rwanda was just the name of any other African country.
+However, in early 2022, the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership.
+[^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible With this agreement, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist 2023).
 Brexit.
 The UK’s departure from the European Union has been controversial since its inception.
-Even eight years post-exit, debates persist regarding the soundness of the decision.
-The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega, 2018).
-Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving, 2019).
-Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022).
-Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
+The division among the population was already evident during the tightly contested 2016 referendum(Arnorsson and Zoega 2018).
+Additionally, misinformation and disinformation have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022).
+Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from Eastern Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
 NHS.
-Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the health care system (The Guardian, 2022; The Independent, 2024).
-As a result, the NHS has become a hot topic of discussion among the UK population.
-Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick et al., 2015).
+Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the healthcare system (The Guardian 2022; The Independent 2024).
+Public opinion is divided, with some supporting that the NHS receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015).
 This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
@@
 Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years.
-Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing.
-Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi, 2021).
-Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021).
-Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
-Threads Comments Year
-Rwanda 0.05 1.87 2020, 2022
-Brexit 0.37 11.57 2020 - 2022
-NHS 0.34 6.20 2020 - 2022
-Vaccine 1.42 28.11 2019, 2022
-Feminism 0.94 17.79 2020 - 2022
-EU 0.2 12.4 2020 - 2022
-Red Cross 0.02 0.2 2020 - 2022
-SubSaharan 0.08 1.5 2020, 2022
-Surgery 0.5 3.9 2019, 2022
-Woman 2.1 19.3 2020 - 2022
-Table 2. Numbers, in millions, of threads, comments, and time period for each trigger word.
+Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021).
+Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al.
+2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
 Feminism.
-Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz, 2019).
-Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al., 2022).
-[^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women.
-This has led to the formation of online echo chambers (Farrell et al., 2019) where such views are amplified and propagated.
-Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
+Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019).
+Some social media influencers have built a large audience by sharing and promoting harmful content against women.
+This has led to the formation of online echo chambers (Farrell et al.
+2019) where such views are amplified and propagated.
+Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope, and unreasonable behavioural demands.
 Keyword Subreddits
@@
 CoronavirusUK, Coronavirus, wuhan_flu
-Table 3. Subreddits selected for each trigger word.
+Table 2: Subreddits selected for each trigger word.
 Feminist Rwanda Brexit NHS Vaccine
@@
 Space 49 51 4.1 (3.8, 4.5) 50 50 9.7 (8.2, 11.3) 49 51 6.3 (6.1, 6.6) 49 51 6.1 (5.7, 6.4) 49 51 4.1 (3.5, 3.9)
-Semantic 48 52 4.6 (3.9, 5.2) 50 50 9.8 (6.6, 14.9) 47 53 3.2 (2.6, 3.7) 53 47 13.4 (8.7, 18.1) 49 51 5.0 (3.1, 6.9)
+Contextual 48 52 4.6 (3.9, 5.2) 50 50 9.8 (6.6, 14.9) 47 53 3.2 (2.6, 3.7) 53 47 13.4 (8.7, 18.1) 49 51 5.0 (3.1, 6.9)
 Temporal - 49 51 7.2 (2.9, 11.6) - - 50 50 5.9 (4.1, 7.6)
-Table 4.
-Percentage of comments per thread before and after each trigger appearance (Hypothesis 1).
-Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons.
+Table 3: Percentage of comments per thread before (Be) and after (Af) each trigger appearance (Hypothesis 1).
+Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Contextual, comparisons.
 Bold scores indicate statistically significant results.
 Detailed distribution of comments can be seen in Appendix B, see Figure 8.
-3.2. Data collection
-Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al., 2020).
-We searched for exact matches of specific keywords in comments from 2019 to 2022, and then collected the entire discussion thread for each extracted comment [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume..
-We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al., 2017).
-Posts marked as “[deleted]” by the API were also removed.
-Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
-As a final filtering step we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word.
-Previous research has shown that Reddit posts typically receive the majority of interactions within a short window after the post is created (Thukral et al., 2018).
-Moreover, early responses are often important in shaping the direction of the conversation (Singer et al., 2016; Horne et al., 2017).
-Considering the fast-paced nature of Reddit, the 30 minutes window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data.
-Table 2 shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered.
-3.3. Research design
-We identify the causal effect of trigger points with the difference-in-difference estimator (DiD) which systematically compares a treatment with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994).
+3.2 Data collection
+Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al.
+2020).
+We searched for exact matches of specific keywords in comments from 2019 to 2022 and then collected the entire discussion thread for each extracted comment.
+[^2] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.
+We excluded comments in languages other than English, identified as such using a fastText-based language identifier (Bojanowski et al.
+2017).
+Comments marked as “[deleted]” by the API were also removed.
+Following this, only threads that contained comments before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
+As a final filtering step, we limited our analysis to comments 30 minutes before and 30 minutes after the first mention of the triggering word.
+Previous research has shown that Reddit comments typically receive the majority of interactions within a short window after the comment is created (Thukral et al.
+2018).
+Moreover, early responses are often important in shaping the direction of the conversation (Singer et al.
+2016; Horne, Adali, and Sikdar 2017).
+Considering the fast-paced nature of Reddit, the 30-minute window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data [^3] Table 7 in Appendix B shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered..
+3.3 Research design
+We identify the causal effect of trigger points with the difference-in-difference estimator (DiD), which systematically compares a treatment with a control group (Angrist and Pischke 2009; Card and Krueger 1994).
 It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group.
@@
 There should be a more substantive behavioural change for the treatment group than for the control group.
-3.3.1. Control groups
+Control groups
 We systematically compare treatment and control groups along three different dimensions.
@@
 Discussions on Reddit are structured in different fora, also called subreddits.
-Aiming to investigate the different effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 3).
+Aiming to investigate the effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 2).
 We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
-2. Semantics
-We also compare trigger words with a term that shares a similar meaning, but that does not go against the deeper underlying transgressions responsible for trigger responses.
+2. Contextual
+We also compare trigger words with terms that appear in a similar context, but that does not go against the more profound underlying transgressions responsible for trigger responses in similar contexts.
 To members of the subreddits in Table 2, using an explicit trigger word or its more harmless control term should make a real difference.
@@
 to Rwanda.
-3. Time
+Treatment and control groups are from the same subreddits in the same year.
+3. Temporal
 In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time.
-For “Rwanda” we compare discussions in subreddits likely to incite reactions in 2020 to discussions in 2022.
+For “Rwanda”, we compare discussions in subreddits likely to spark reactions in 2020 to discussions in 2022.
 For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic).
 Again, we expect that users from the subreddits in Table 2 are triggered by those only in 2022 but not in the respective earlier years.
-3.3.2. Statistical analysis
+Treatment and control groups share the same subreddits but are from different years.
+Statistical analysis
 We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 3) and 48 experiments for H2 (Table 5).
 For each experiment, we prepare the data in the same way.
-The treatment/placebo is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread.
-To test H1, we count the number of posts for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group.
-We then calculate the share of posts before and after the intervention.
-For example, if there are 8 posts during the 30 minutes before the trigger word and 12 posts during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention.
-For experiments with H2, we count the number of the respective animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
-We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke, 2009):
+The treatment is the first mention of a trigger word in a discussion thread, i.e., there is only one well-defined treatment per thread.
+To test H1, we count the number of comments for each discussion thread 30 minutes before and 30 minutes after the treatment in the treatment group and the placebo treatment in the control group.
+We then calculate the share of comments before and after the intervention.
+For example, if there are 8 comments during the 30 minutes before the trigger word and 12 comments during the 30 minutes after the trigger word, we normalize the data of this discussion thread to 40% before intervention and 60% after intervention.
+For experiments with H2, we count the number of comments expressing animosity in a thread and then normalize the data as a share of animosity per thread before and after the trigger word.
+We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke 2009):
 […]
 where proportion is the per-thread-normalised target feature and the interaction term is given by
-$\text{interaction}=\text{trigger\_dummy}\times\text{treatment\_dummy}.interaction = trigger_dummy × treatment_dummy$ .
-Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al., 2023).
+[…]
+Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al.
+2023).
 It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time.
 We investigate the plausibility of the assumption in section 4.3.
-3.4. Features
+3.4 Features
 We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions.
-We simply use post counts for the first hypothesis (H1) related to user engagement.
+We use comment counts for the first hypothesis (H1) about user engagement.
 The second hypothesis (H2) requires a measure of animosity.
-To identify relevant features that can operationalise animosity, we build on various language models that follow the RoBERTa architecture (Liu et al., 2019) and are trained on social media corpora, specifically X (Twitter).
-Even though posts on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text.
-Hence, we expect the trained models to perform well in either context (Priya et al., 2019).
-3.4.1. User engagement (Hypothesis 1)
+To identify relevant features that can operationalize animosity, we build on various language models that follow the RoBERTa architecture (Liu et al.
+2019) and are trained in social media corpora, specifically X (Twitter).
+Even though comments on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text.
+Hence, we expect the trained models to perform well in either context (Priya et al.
+2019).
+User engagement (Hypothesis 1)
 Our first hypothesis suggests that trigger words lead to higher user engagement.
-We simply rely on the number of posts to investigate this.
-That is, we expect to see an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
-3.4.2. Animosity (Hypothesis 2)
-Hypothesis 2 suggests that trigger words cause more animosity in a debate.
-We operationalise this with controversial comments, the presence of negative sentiment, anger, and different forms of hate speech.
+We rely on the number of comments to investigate this and expect an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
+Animosity (Hypothesis 2)
+Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalize this with controversial comments, negative sentiment, anger, and different forms of hate speech.
 Controversiality
@@
 We expect trigger words to cause more controversiality.
-Figure 2. Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
+Figure 2: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
 Sentiment analysis
-For sentiment analysis, we rely on the model twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) which identifies a post as negative, neutral, or positive.
-This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019).
-The sentiment in social media content has often been used to gain insights into the information related to controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender biases (Marjanovic et al., 2022).
-We expect trigger words to cause more negative sentiments.
+For sentiment analysis, we use twitter-roberta-base-sentiment-latest (Loureiro et al.
+2022) which identifies a comment as negative, neutral, or positive.
+This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019).
+As we expect trigger words to cause more negative sentiments, we use sentiment analysis to investigate this, similarly to work using it to gain insights into how controversial topics such as vaccination (Melton et al.
+2021), politics (Guimaraes et al.
+2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022) are discussed online.
 Negative emotions
-We rely on twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022) to assign one or more emotions to each tweet.
-This model is trained with data from the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018), covering 11 different emotions [^6] Detailed list of emotions can be found in Appendix A.1).
-For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
-Figure 3. Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
+We leverage twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al.
+2022) to assign one or more emotions to each tweet.
+This model is trained on the Semeval 2018 Affect in Tweets task (Mohammad et al.
+2018), covering 11 different emotions.
+[^6] Detailed list of emotions can be found in Appendix A.1) For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
+Figure 3: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
 Hate speech
-Finally, we use the twitter-roberta-base-hate-multiclass model (Antypas and Camacho-Collados, 2023) to detect hate speech.
-It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups [^7] Detailed list of categories can be found in Appendix A.2).
-The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al., 2021), women’s rights (Farrell et al., 2019), and health issues (Walter et al., 2022).
-Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist, racism for Rwanda, while we consider all hateful comments jointly for NHS, Brexit and Vaccine.
-4. Results
-In this section, we present the results for the two hypotheses tested in this paper.
-4.1. User engagement (Hypothesis 1)
-Our initial hypothesis is that if trigger words significantly influence online discussions, there should be more messages after mentioning these terms.
+Finally, we use the twitter-roberta-base-hate-multi-class model (Antypas and Camacho-Collados 2023) to detect hate speech.
+It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups.
+[^7] Detailed list of categories can be found in Appendix A.2 ) The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al.
+2021), women’s rights (Farrell et al.
+2019), and health issues (Walter et al.
+2022).
+Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist and racism for Rwanda.
+For NHS, Brexit and Vaccine we consider all hateful comments jointly.
+3.5 Validation of Annotation Model
+The primary aim of our research is to use simple and computationally efficient models that reliably extract signals relevant to our hypotheses.
+To confirm the validity and robustness of our findings, we implemented an additional validation procedure to ensure consistency across the classifiers employed in our analysis.
+We sampled 100 Reddit comments from our data collection, with 20 comments selected from the subreddits associated with each trigger point (Table 2).
+The collected comments were independently annotated by three experienced NLP and CSS researchers for sentiment, emotion, and the presence of hate speech.
+[^8] We include the annotation guidelines in Appendix A.4.
+Each comment was assigned a label based on the majority vote, while disagreements were resolved through discussion.
+When considering the difficulty of the annotation task, the annotations present above-average agreement scores for sentiment and hate speech, with Fleiss’ $\kappa$ scores of 0.52 and 0.56, respectively.
+For the multi-label emotion annotation task, where annotators had to choose from a set of 12 labels,the Krippendorff’s alpha score (Krippendorff 2011) reached 0.24, which is in line with other similar multi-label tasks with a high number of labels (Mohammad et al.
+2018; Antypas et al.
+2022; Ousidhoum et al.
+2019).
+We fine-tune the base version of BERT-cased, XLNet, and RoBERTa for sentiment analysis, emotion recognition, and hate speech detection on the same datasets as the models we used for our experiments, as explained in Section 3.4 (Rosenthal, Farra, and Nakov 2017; Mohammad et al.
+2018; Antypas and Camacho-Collados 2023) to ensure fairness.
+We further compare our results to those of five pre-trained models publicly available on HuggingFace.
+Specifically, two sentiment classifiers trained on Reddit data999 reddit-dBERT: https://huggingface.co/akshataupadhye/finetuning-sentiment-model-reddit-data, reddit-XLNET: https://huggingface.co/minh21/XLNet-Reddit-Sentiment-Analysis , hateBERT (Caselli et al.
+2021) and one emotion classifier [^10] BERT-emotion: https://huggingface.co/ayoubkirouane/BERT-Emotions-Classifier.
+All models are evaluated in the aforementioned manually labelled dataset.
+Table 4 presents the average micro F1-scores achieved by the models across each task.
+We opted for the micro F1-score as it provides a more realistic representation of the results, given the limited sample and heavily imbalanced nature of our dataset.
+[^11] Detailed dataset distribution is available in Appendix A.3 The Twitter-RoBERTa models achieved the highest scores on all tasks.These scores are matched by XLNet-base for sentiment analysis and XLNet for emotion detection, and the fine-tuned models RoBERTa and XLNet for hate speech detection.
+Overall, the tested models performed comparably well, except for two off-the-shelf sentiment classifiers that achieved scores below 50%.
+Sentiment Emotion Hate
+t-RoBERTa 0.76 0.50 0.88
+RoBERTa 0.75 0.47 0.88
+XLNet 0.76 0.45 0.88
+BERT 0.74 0.40 0.85
+reddit-dBERT 0.45 - -
+reddit-XLNet 0.31 - -
+BERT-emotion - 0.45 -
+hateBERT - - 0.80
+Random 0.30 0.20 0.07
+Table 4: Average micro-F1 scores of the models tested for the sentiment analysis, emotion, and hate speech detection tasks. Random indicates a naive baseline that randomly selects labels.
+4 Results
+4.1 User engagement (Hypothesis 1)
+According to our first hypothesis, trigger points cause more comments.
 The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 3).
-In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all the terms we selected.
-It is largest with Rwanda where we observe a 10%-points (Table 4) increase.
+In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all terms.
+It is the largest with Rwanda where we observe a 10%-points (Table 3) increase.
 To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group.
 In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate.
-The lowest value is a 3.2%-point increase for the semantic comparison in Brexit.
-The largest increase is for NHS in the semantic control group, where we observe a 13.4% point increase.
+The lowest value is a 3.2%-point increase for the contextual comparison in Brexit.
+The largest increase is for NHS in the contextual control group, where we observe a 13.4% point increase.
 When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent.
@@
 This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial.
-Similar patterns are also observed in the semantic and temporal control groups for Rwanda (Appendix B, Figure 9).
-Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our initial hypothesis on the effect of trigger words on online discussions.
+Similar patterns are also observed in the contextual and temporal control groups for Rwanda (Appendix B, Figure 10).
+Overall, our analysis of the influence of trigger words across different control groups indicates a clear increase in engagement for each trigger word analysed, confirming our first hypothesis.
 Trigger Setting Feature DiD CI
@@
 Hate -0.1 (-0.5, 0.4)
-Semantics Contr 1.4 (0.7, 2.1)
+Context. Contr 1.4 (0.7, 2.1)
 N.Sent 2.6 (2.0, 3.1)
@@
 Sexism 0.6 (0.1, 1.1)
-Semantics Contr 0.1 (-0.8, 1.1)
+Context. Contr 0.1 (-0.8, 1.1)
 N.Sent 3.0 (2.3, 3.6)
@@
 Hate -0.3 (-0.9, 0.3)
-Semantics Contr 4.5 (-3.2, 12.2)
+Context. Contr 4.5 (-3.2, 12.2)
 N.Sent 3.5 (-1.2, 8.3)
@@
 Racism 1.8 (0.0, 3.6)
-Semantics Contr 6.6 (0.2, 13.0)
+Context. Contr 6.6 (0.2, 13.0)
 N.Sent 5.5 (1.5, 9.6)
@@
 Hate 0.4 (0.0, 0.8)
-Semantics Contr 2.2 (1.1, 3.4)
+Context. Contr 2.2 (1.1, 3.4)
 N.Sent 5.4 (4.7, 6.1)
@@
 Hate -2.0 (-5.3, 1.3)
-Table 5. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
-4.2. Animosity (Hypothesis 2)
-Our second hypothesis suggests that trigger words may increase the frequency of polarising messages and induce animosity in the discussion.
-Figure 4. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
+Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
+4.2 Animosity (Hypothesis 2)
+Our second hypothesis suggests that trigger words may increase the frequency of polarised messages and lead to more animosity in the discussion.
+[^12] An example of such a discussion is presented in Appendix E.
+Figure 4: Average controversiality per thread for the treatment and control (space) groups of Rwanda.
 Controversiality.
-We analyse the average controversiality of each thread and compare the proportion of controversial threads before and after the introduction of the trigger words.
+We analyse the average controversiality of each thread and compare the proportion of controversial comments before and after the introduction of the trigger words.
 Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word.
-Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the semantic comparison for Rwanda—a causal effect of 6.6% points.
+Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the contextual comparison for Rwanda—a causal effect of 6.6% points.
 However, significant differences in controversiality do not appear across all settings.
-Notably, there seems to be no increase in controversial comments when considering Feminism, and in the temporal and semantic settings for Vaccine and NHS, respectively.
+Notably, there seems to be no increase in controversial comments for Feminism, and in the temporal and contextual settings for Vaccine and NHS, respectively.
 Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence.
 The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points.
-A similar trend is seen in both the semantic and temporal control settings, as documented in Appendix B, Figure 8).
+The contextual and temporal control settings follow a similar trend, as documented in Appendix B, Figure 9.
 (a) Rwanda
 (b) Feminism
-Figure 5. Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
+Figure 5: Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
 Sentiment & emotion.
@@
 In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
-When considering the negatively charged comments of the treatment and control groups for each trigger word, Table 5 indicates a clear trend.
-We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the semantics comparison for NHS.
-The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and semantics experiments.
+Table 5 indicates a clear trend for the negatively charged comments of the treatment and control groups for each trigger word.
+We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the contextual comparison for NHS.
+The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and contextual experiments.
 Let us consider the vaccine trigger, Figure 3, as an example.
@@
 Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned.
-However, the treatment group shows more messages than the control group, suggesting that the trigger words affects the expression of negative sentiment.
+However, the treatment group displays a noticeably higher volume of messages around the trigger term, suggesting that triggers amplify the expression of negative sentiment compared to the control condition.
 Similarly to the increase in negative sentiment, trigger words systematically cause more anger.
 We observe a positive causal effect in all conditions, with only 3 non-significant effects.
-Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine semantic comparison).
-As another visual example, the overall number of posts expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure LABEL:fig:emotion_space_vaccine.
+Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine contextual comparison).
+As another example, the overall number of comments expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 12, Appendix B.
 This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
 Hate speech.
-As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of hate speech or abusive language for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism.
+As a final indicator of animosity, we also consider hate speech and abusive language, more specifically any kind of toxicity for Brexit, NHS, and Vaccine, and the more domain-specific racism-related hate speech for Rwanda and sexism-related hate speech for Feminism.
 As seen from Table 5, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment.
-We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and semantic comparison) and Rwanda (space group).
-We illustrate our findings again in a visualisation, where we consider all messages irrespective of the thread they are posted in.
+We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and contextual comparison) and Rwanda (space group).
+We illustrate our findings in a visualization, where we consider all messages irrespective of the thread in which they are posted.
 We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger.
 Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
-4.3. Testing the Parallel Trends Assumption.
+4.3 Robustness of Results
+We conduct a series of robustness checks designed to test the sensitivity of our results to assumptions and alternative specifications.
+Testing the Parallel Trends Assumption.
 The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time.
@@
 The first argument is a theoretical one.
-In neither of our experimental designs, we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word.
-Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (semantics), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
+In neither of our experimental designs did we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word.
+Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (contextual), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
 We implement event studies for each experiment to further empirically support our argument.
 The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention.
-Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel.
-If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al., 2023).
-Figure 6. Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
-For each of our overall 60 experiments, we divide the 30 minutes before and after the first mention of a trigger word into six distinct time intervals $\{[-30,-20[,[-20,-10[,[-10,0[,[0,10[,[10,20[,[20,30[\}${ [ - 30 , - 20 [ , [ - 20 , - 10 [ , [ - 10 , 0 [ , [ 0 , 10 [ , [ 10 , 20 [ , [ 20 , 30 [ } instead of just two.
-This allows us to follow, in particular, the pre-treatment trends at a higher granularity.
-We calculate each discussion thread’s share of (hostile) messages in the six intervals.
-Averaging over all these data points per period, we subtract the control group from the treatment group and calculate the difference of each observation from the first observation period.
+Plotting the treatment effects at different time points before the treatment, the event study helps visualize whether the pre-treatment trends in the treatment and control groups are indeed parallel.
+If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al.
+2023).
+[…]
 Figure 7 charts the time-wise means over all 60 event studies.
@@
 [^13] See Appendix for figures that include individual event studies for all experiments from H1 and H2.
-5. Conclusion
-In this paper, we present the concept of trigger words and study how they affect social media conversations in UK politics.
-We report empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine).
-In short, our results show that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
-Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word.
-We consistently observe more posts across all trigger words and control settings.
+Placebo Treatment.
+To further validate our Difference-in-Differences (DiD) estimates, we implement a placebo treatment analysis.
+The core logic of this approach is to check whether the estimated treatment effects could be driven by other trends earlier or later than the actual intervention.
+In the placebo analysis, we systematically shift the assumed treatment time in 10-minute intervals, ranging from 30 minutes before to 30 minutes after the actual intervention, i.e., the comment with the first mention of the trigger word in a discussion thread.
+At each placebo point, we re-estimate all DiD models as if the treatment had occurred at that alternative time.
+If our results capture a genuine causal effect, we expect a pronounced treatment effect at the actual intervention time ($t=0$), and near-zero or statistically insignificant estimates at all placebo points before and after.
+Figure 6 visualizes these placebo tests by plotting the precision-weighted average treatment effects and their confidence intervals across all experiments measuring animosity (see Table 5).
+At the actual treatment point $t=0$, which compares message content in the 30 minutes before and after the first appearance of the trigger word ([30,0[[-30,0[[ - 30 , 0 [ vs.
+$[0,30]$), we observe a clear and substantial effect (1.99).
+We then shift the treatment window rightward in 10-minute steps—first comparing [20,10[[-20,10[[ - 20 , 10 [ to $[10,40]$, and so on—until the actual intervention fully exits the analysis window at $t=30$.
+This generates a steady decline in estimated effects, which converge to zero, as expected if no real treatment occurred at these placebo times.
+Similarly, shifting the window to earlier times (e.g., comparing [60,30[[-60,-30[[ - 60 , - 30 [ to $[-30,0]$) reveals a decreasing trend in effect size, though some residual difference remains.
+In conclusion, given the clear peak of the effect at the timing $t=0$, we are confident in the robustness of our approach to defining the moment of the actual treatment moment.
+Figure 6: Placebo treatments every 10 minutes, ranging from 30 minutes before to 30 minutes after the actual treatment point.
+Precision weighted average treatment effects and 95% confidence intervals across all experiments.
+Figure 7: Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
+5 Conclusion
+This paper establishes trigger words as a concept in online debates.
+Since our work represents the first systematic computational study of this concept, our research design and case selection are geared to clearly identify such behaviour in the most likely cases.
+To do so, we performed a large-scale analysis about UK politics on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS, and Vaccine).
+The results show that trigger words cause more user engagement, and this activity is marked by higher levels of animosity.
+This confirms the first hypothesis (H1), which suggests more user interaction after the appearance of a trigger word.
 Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects.
@@
 We also investigate the causal effect of trigger words on hate speech.
-Where we can, we rely on a model that is capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—.
+Where we can, we rely on a model capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—.
 Our findings similarly confirm that trigger words play a causal role in this aspect as well.
-6. Limitations
-We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries).
+From this work, we can certify that trigger points exist in online communication, and we have provided a framework to measure the extent of them in terms of user engagement and animosity.
+Future work will analyse this phenomenon more comprehensively.
+We aim to broaden the scope, study how prevalent trigger points are in different countries and languages, and seek to better understand further behavioural consequences from trigger points.
+It is our intention to evaluate the extent to which trigger points are responsible for hateful online communication.
+We also strive to automate the trigger point detection process to detect the respective words in online conversations when we do not know them ex-ante.
+6 Limitations
+We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset ($>$100 million entries).
 Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments.
-Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al., 2023).
-Selecting words as trigger points and their corresponding subreddits offers room for improvement.
+Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al.
+2023).
+Our study does not attempt to map how pervasive trigger points are in online debates.
 The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community.
-Similarly, we only investigate a very select set of potential trigger words in the context of the UK.
-However, our study had different goal: We consider our work to be a pilot study that intends to investigate whether the trigger points exist in online settings at all.
-Based on the empirical evidence we find, we can be confident to have found relevant empirical evidence that warrants further investigation.
-7. Ethics Statement
-All data utilised in our research is sourced from publicly available information or collected using the official Reddit API.
-All information is provided in an aggregated fashion without reporting sensitive information from individual users.
-Appendix A Model Categories
-A.1. Emotion Categories
+Similarly, we acknowledge the limitations and potential biases of our work, as we only investigate a small set of potential trigger words in the context of the UK and limit our data to comments written in English in a limited selection of subreddits.
+The choice of trigger words is informed by a contextual analysis of the UK politics, and the Reddit threads have been selected due to the often polarised conversations they host around such topics.
+However, this is the first study that analyses trigger points’ concept validity in online conversations.
+Based on our empirical evidence, we can be confident that we have found relevant evidence that warrants further investigation.
+7 Ethics Statement
+All data used in our research is sourced from publicly available information or collected using the official Reddit API.
+It is provided in an aggregated way and without reporting sensitive information from individual users.
+The automatic methods used can be abused by those with the power to repress people with opposing opinions.
+Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use.
+Researchers interested in replicating our experiments can contact us directly.
+Appendix A Model Selection
+A.1 Emotion Categories
 The twitter-roberta-base-emotion-multilabel model classifies each entry in one or more of the following classes: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust.
-A.2. Hate Speech Categories.
+A.2 Hate Speech Categories.
 The twitter-roberta-base-hate-multiclass model classifies each entry in one of the following classes: not_hate, sexism, racism, religion, other, sexual_orientation, disability
+A.3 Validation Setting
+Table 6 displays the distribution of comments in regards to emotion, sentiment and hate speech for the manually annotated dataset collected.
+Emotion Hate Speech Sentiment
+no emotion 29 not hate 88 negative 56
+anger 26 sexism 7 neutral 37
+disgust 23 racism 3 positive 7
+pessimism 19 other 2
+anticipation 15
+sadness 9
+surprise 5
+optimism 5
+joy 2
+Table 6: Distribution of comments regarding emotion, sentiment, and hate speech categories on the manually annotated dataset.
+A.4 Annotation Guidelines
+Below we present the guidelines presented to annotators for the Sentiment Analysis, Emtion classification, and Hate Speech detection tasks.
+Sentiment Analysis
+We will show you Reddit comments that can be positive, negative, or neutral based on their sentiment.
+Sentiment Categories
+  • Positive comments should be labelled: Positive
+  • Neutral comments should be labelled: Neutral
+  • Negative comments should be labelled: Negative
+Examples
+  • “‘Just finished my morning workout and feeling fantastic! #HealthyLiving #FitnessGoals“‘ is a positive sentence.
+  • “‘Stuck in traffic again! Late for work and already stressed out. #TrafficJam“‘ #MondayBlues is a negative sentence.
+  • Factual sentences such as “‘The party was in front of the house“‘ are neutral sentences.
+Notes: We have also added a comment column where you can share your thoughts, comments, or doubts.
+If you are unsure whether you should annotate the tweet as positive, negative or neutral, you may choose the label that suits it best and a comment about that in the comment column.
+The annotation of an ironic comments depends on the context. While we acknowledge the lack of context, if you think the tweet is mocking someone (or a group of people), you should label it negative.
+Emotion Classification
+You will be presented with Reddit comments and asked to select the option that best describes the text.
+You may choose as many emotions as you believe are present — from 0 to 11 — based on the content and tone of the comment.
+Emotion Categories:
+  • Anger – includes annoyance, rage
+  • Anticipation – includes interest, vigilance
+  • Disgust – includes disinterest, dislike, loathing
+  • Fear – includes apprehension, anxiety, terror
+  • Joy – includes serenity, ecstasy
+  • Love – includes affection
+  • Optimism – includes hopefulness, confidence
+  • Pessimism – includes cynicism, no confidence
+  • Sadness – includes pensiveness, grief
+  • Surprise – includes distraction, amazement
+  • Trust – includes acceptance, liking, admiration
+Notes: If no emotion is clearly present, you may leave all options unchecked.
+Choose all emotions that are clearly expressed or implied in the comment.
+If you’re unsure, use the comment column to share your reasoning or highlight ambiguity.
+Comments may express multiple emotions simultaneously, for example, both sadness and anger.
+Hate Speech Detection
+You will be presented with Reddit comments and asked to identify whether they contain hate speech, and if so, which target group is being attacked or discriminated against.
+Please choose only one of the following categories that best represents the nature of the content:
+Hate Speech Categories:
+  • Sexism: Content that targets or demeans people based on their gender (e.g., women, men, non-binary individuals) or reinforces harmful gender stereotypes.
+Example: “Women should stay in the kitchen where they belong.”
+  • Racism: Content that expresses hostility, prejudice, or dehumanization based on race, ethnicity, or national origin. Example: “All [ethnic group] are criminals and should be deported.”
+  • Disability: Content that mocks, belittles, or degrades people with physical, mental, or developmental disabilities. Example: “You’re so stupid, you must be on the spectrum.”
+  • Sexual Orientation: Hate directed toward someone because of their actual or perceived sexual orientation (e.g., gay, lesbian, bisexual, etc.).
+Example: “Being gay is a disease that should be cured.”
+  • Religion: Disparagement, dehumanization, or threats against individuals or groups based on religious beliefs or identities. Example: “All Muslims are terrorists.”
+  • Other: Use this label for hate speech that targets a group not covered above, such as socioeconomic status, age, appearance, political affiliation, etc.
+Example: “Old people are useless and should just die already.”
+  • Not Hate: Use this label if the comments does not contain hate speech. Comments that are neutral or non-discriminatory.
+Notes: If you are unsure, select the label you think fits best and add a comment explaining your reasoning.
+A single post may mention multiple groups in such a case, select the one that is most explicit or harmful, and explain in the comment column.
 Appendix B Extended Results
+Table 7 presents the total number of comments and threads extracted from Reddit for each keyword, along with the corresponding time periods during which the data was collected.
+Table 8 presents the number of comments included in each experiment for both the treatment and control groups.
 Figure 8 displays the number of comments posted before and after the introduction of each trigger point on the data collected.
-Figure 8 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
-Figure 9 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
+Figure 9 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
+Figure 10 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
 Figure 11 displays the distribution of negative comments, based on their sentiment, posted before and after the introduction of the Brexit, Feminism, Rwanda, and NHS trigger points in the Space control setting.
+Figure 12 displays the distribution of comments conveying anger and those that do not before and after the introduction of the Vaccine trigger point in the Space control setting.
+Threads Comments Year
+Rwanda 0.05 1.87 2020, 2022
+Brexit 0.37 11.57 2020 - 2022
+NHS 0.34 6.20 2020 - 2022
+Vaccine 1.42 28.11 2019, 2022
+Feminism 0.94 17.79 2020 - 2022
+EU 0.2 12.4 2020 - 2022
+Red Cross 0.02 0.2 2020 - 2022
+SubSaharan 0.08 1.5 2020, 2022
+Surgery 0.5 3.9 2019, 2022
+Woman 2.1 19.3 2020 - 2022
+Table 7: Numbers, in millions, of threads, comments, and time period for each trigger word.
+Trigger Experiment Treatment Control
+Feminist space 264,495 16,503,956
+contextual 163,740
+Rwanda space 28,770 668,641
+contextual 2,587
+temp 5,896
+Brexit space 688,600 10,793,036
+contextual 111,524
+vaccine space 348,641 7,685,095
+contextual 12,986
+temp 11,841
+NHS space 509,458 5,275,917
+contextual 1,276
+Table 8: Number of comments present for each trigger in each experiment for both the treatment and control subsets.
 (a) Brexit
@@
 (e) Vaccine
-Figure 7. Absolute count of posts before and after each trigger point.
+Figure 8: Absolute count of comments before and after each trigger point.
 (a) Temporal Setting
-(b) Semantic Setting
-Figure 8. Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
-(a) Semantic Comparison
+(b) Contextual Setting
+Figure 9: Average controversiality per thread in the Temporal and Contextual settings and their respective treatment groups, for the trigger point Rwanda.
+(a) Contextual Comparison
 (b) Temporal Comparison
-Figure 9. Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
+Figure 10: Rwanda Treatment and Control distributions of comments in the Contextual and Temporal settings.
 (a) Brexit
@@
 (d) NHS
-Figure 10. Distribution of negatively charged comments in the Treatment and Control (Space) groups.
+Figure 11: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
+Figure 12: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
 Appendix C Extended Test of Parallel Trends Assumption
+Figures 13 a & b display the results of the Parallel Trends Assumption testing for all experiments as described in H1 and H2.
 (a) Experiments from H1.
 (b) Experiments from H2.
-Figure 11. Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
+Figure 13: Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
 Appendix D Computational Resources
 An NVIDIA GeForce RTX 4090 GPU was utilised for the experiments conducted. 170 hours spent for the inference process of sentiment analysis, emotion detection, and hate speech detection models.
+Appendix E Selected Reddit Comments
+Below we present a Reddit discussion and how it evolves when the trigger ”feminism” appears. The comment containing the trigger is bolded.
+WARNING: the following comments contain sensitive and potentially offensive language. Reader discretion is advised.
+”This image though…”
+”Haunting for us, isn’t it? But not for the females.”
+”Well I’m sure most females would find this horrifying too.”
+”You a female? Why are you on this Sub?”
+”Because I’m against feminism?”
+”Well, it’s better if these Posts are seen by all the Men out there. Get the taste of true Chameleon female nature.”
+”Too late. I already saw it.”
+”There are many women like us on this sub .”
+”Yeah I’m a woman and I don’t like it. Women are not a hive mind.”
+”Women are not a hive mind. You should go out and talk to one maybe.”
+”My fucking eyes. Sometimes I wish I was blind.”
+”When Male dies, it’s ””sometimes”” sad; when Female dies………….CRIMEEEEE!!!!
+”The fucked up world we live in”

### v4

@@
-Words as Trigger Points in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit
-Abstract
-Political debates on social media sometimes flare up.
-From that moment on, users engage much more with one another; their communication is also more emotional and polarised.
-While it has been difficult to grasp such moments with computational methods, we suggest that trigger points are a useful concept to understand and ultimately model such behaviour.
-Established in qualitative focus group interviews to understand political polarisation (Mau, Lux, and Westheuser 2023), trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned.
-In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned.
-Our paper finds that these trigger points also exist in online debates.
-We examine online deliberations on Reddit between 2020 and 2022 and collect $>$100 million comments from subreddits related to a set of words identified as trigger points in UK politics.
-Analysing the comments, we find that trigger words increase user engagement and animosity, i.e., more negativity, hate speech, and controversial comments.
-Introducing trigger points to computational studies of online communication, our findings are relevant to researchers interested in affective computing, online deliberation, and how citizens debate politics and society in light of affective polarisation.
-1 Introduction
-Online debates may suddenly escalate, and users intensely engage in emotional and confrontational deliberations.
-While these turning points are difficult to model and predict, they are essential moments in online communication.
-They mark the moment when a discussion is likely to get off the rails, and social media users stop debating with one another and instead debate against one another.
-Our study introduces trigger points as a useful concept to understand why and when social media users engage in such behaviour.
-Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions.
-They open the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
-Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
-Trigger points are related to how individuals perceive themselves in society.
-People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions.
-These deeply rooted beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world.
-People feel triggered if these normative understandings are questioned.
-More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract.
-Trigger points jeopardise what individuals consider the fabric of society and their own role in it.” The authors suggest that trigger points can affect individuals through four mechanisms.
+Causal Effects of Trigger Words in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit
+Abstract.
+Political debates on social media often escalate quickly, leading to increased engagement as well as more emotional and polarised exchanges.
+Trigger points (Mau et al., 2023) represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is being questioned.
+Analysing Reddit discussions, we examine how trigger points shape online debates and assess their impact on engagement and affect.
+Our analysis is based on over 100 million comments from subreddits centred on a predefined set of terms identified as trigger words in UK politics.
+We find that mentions of these terms are associated with higher engagement and increased animosity, including more controversial, negative, angry, and hateful responses.
+These results position trigger words as a useful concept for modelling and analysing online polarisation.
+Affective polarization, Causal effect, Computational social science, Hate speech, Reddit, Trigger words, UK politics
+* Equal contribution
+journalyear: 2026copyright: ccconference: 18th ACM Web Science Conference; May 26–29, 2026; Braunschweig, Germanybooktitle: 18th ACM Web Science Conference (WebSci ’26), May 26–29, 2026, Braunschweig, Germanydoi: 10.1145/3795766.3799750isbn: 979-8-4007-2504-3/2026/05ccs: Applied computing Sociologyccs: Information systems Social networksccs: Computing methodologies Discourse, dialogue and pragmatics
+1. Introduction
+Online debates may suddenly escalate, prompting users to engage in emotional and confrontational deliberations.
+While this exact moment is challenging to model and predict, it is essential in online communication.
+It signals when discussions derail and users shift from debating to attacking one another.
+Our study introduces trigger words as a useful concept for understanding why and when social media users engage in such behaviour—which is central to accounting for hate speech and other abusive online interactions.
+Figure 1. Consequences of trigger words. Hypothesis 1: Trigger words cause more engagement; Hypothesis 2: Trigger words cause more animosity.
+Trigger words are related to how individuals perceive themselves in society.
+People hold internal models of what counts as acceptable behaviour.
+These models include explicit, rule-like knowledge, e.g., formal norms or laws.
+More importantly, they also contain implicit, hard-to-articulate expectations about which actions are appropriate in a given context.
+Such deeply rooted beliefs about what is fair and what is unfair constitute a moral compass fundamental to navigating the social world.
+People feel triggered if these normative understandings are questioned (Figure 1).
+More formally, Mau et al.
+(2023) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract.
+Trigger points jeopardise what individuals consider the fabric of society and their own role in it.” In our study, we investigate to what extent specific words in online discussions can behave as such trigger points.
+Trigger points can affect individuals through four mechanisms (Mau et al., 2023):
 Inequality.
@@
 Either similar people are treated unequally, or, to the contrary, people who should be treated differently are treated the same.
-For example, social benefits might be at eye level with minimal wage earners.
+For example, social benefits might be at eye level for minimum-wage earners.
 Norm transgression. It is the transgression of what is considered “normal” behaviour. A typical example relates to the lavish lifestyle of the privileged.
 Slippery slope.
-There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it.
-Claims such as “Opening the borders so that a country could be flooded with immigrants” is an example.
+People may fear a society developing norms about appropriate behaviour that go in subjectively undesired directions.
+They may feel they lack the power to do anything about it.
+Claims such as “Opening the borders so that a country could be flooded with immigrants” are an example.
 Behavioural demands.
-When society may be perceived as demanding unreasonable behaviour.
-The role of pronouns is a point in case where the generic masculine pronouns “he/him” might trigger some, and the neutral pronoun “they/them” might trigger others.
-Going against deeply held beliefs about individuals and their societal role causes discomfort.
-People respond to such triggers in an affective behavioural mode, increasing a debate’s intensity and emotionality.
-In the social media context, we expect that this has two direct consequences (Figure 1). Trigger points lead to:
-  • Hypothesis 1: more user engagement, which leads to more messages or comments.
-  • Hypothesis 2: higher levels of animosity, which causes polarisation, negativity, anger, and hate.
-Our study seeks to establish the concept of trigger points in online communication.
-The research design and case selection closely follow this logic and examine whether and how individual words can act as trigger points in online debates.
-To this end, we conduct a large-scale case study and select five words that are ex-ante likely to work as trigger points in the context of UK politics.
-Specifically, we look into the words Rwanda, Brexit, NHS, Vaccine, Feminism.
-Collecting $>$100 million related Reddit comments, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses.
-The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
-  1. 1.
-    Space: specific subreddits where the selected words trigger users.
-  2. 2.
-    Contextual: words that appear in a similar context to the target words but without the triggering component.
-  3. 3.
-    Time: periods identified as relevant for each specific trigger word.
-We find that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
-2 Related Work
-Existing work studies how particular political words or phrases lead to more engagement (Brady et al.
-2017; Hameleers et al.
-2021) and posts with more emotion and animosity (Friedman et al.
-2004; Martin et al.
-2021; Schroth, Bain-Chekal, and Caldwell 2005; Rho and Mazmanian 2020) in various online settings.
+A society may be perceived as demanding unreasonable behaviour.
+The role of pronouns is a case in point where the generic masculine pronouns “he/him” might trigger some, and the neutral pronoun “they/them” might trigger others.
+Because these trigger points cause emotional discomfort, people respond in an affective-behavioural mode, increasing the debate’s intensity and emotionality.
+In the social media context, this has two direct consequences (Figure 1).
+Trigger words lead to: (i) Hypothesis 1: more user engagement, which leads to more messages or comments; and (ii) Hypothesis 2: higher levels of animosity, which causes polarisation, negativity, anger, and hate.
+The goal of our study is to highlight and establish the concept of trigger words in scholarly work on online communication.
+Our research design and case selection closely follow this logic: We want to show that even the mention of a single trigger word can escalate online debates.
+To this end, we select five words that are ex-ante likely to work as trigger words in the context of UK politics (Rwanda, Brexit, NHS, Vaccine, Feminism).
+Collecting >>100 million in related Reddit comments, we utilise text analytics and classifiers to determine whether these words elicit potentially negative, emotional, or harmful responses.
+The analysis builds on a systematic comparison between a treatment and control group in three different scenarios: (1) Space: specific subreddits where the selected words are likely to trigger users; (2) Contextual: words that appear in a similar context to the trigger words but without the triggering component; and (3) Temporal: periods identified as relevant for each specific trigger word.
+2. Related Work
+Existing work studies how particular political words or phrases lead to more engagement (Brady et al., 2017; Hameleers et al., 2021) and posts with more emotion and animosity (Friedman et al., 2004; Martin et al., 2021; Schroth et al., 2005; Rho and Mazmanian, 2020) in various online settings.
 However, they lack a joint conceptual and theoretical framework—something we aim to provide.
-Trigger points are fundamental to understanding polarisation and division in online discussions.
-Given that they are at the root of different conflictive behaviours later on, they help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and their escalation (De Kock and Vlachos 2022; van der Meer et al.
-2023) triggering explicit and implicit ad hominem attacks, hate speech, abusive or offensive language on social media (Talat and Hovy 2016; Davidson et al.
-2017; Basile et al.
-2019; Ousidhoum et al.
-2019; Sap et al.
-2019; ElSherief et al.
-2021; Ocampo et al.
-2023).
-Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al.
-2019), Blodgett et al.
+Trigger words are fundamental to understanding polarisation and division in online discussions.
+Given that they are at the heart of different conflictive behaviours later on, they help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023) triggering explicit and implicit ad hominem attacks, hate speech, abusive or offensive language on social media (Talat and Hovy, 2016; Davidson et al., 2017; Basile et al., 2019; Ousidhoum et al., 2019).Similarly, given a broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al.
 (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model.
-Studying trigger points helps track the origins of bias, assess whether a social media post can lead to toxic comments (Dahiya et al.
-2021), and how language affects polarisation on social media (Karjus and Cuskley 2024).
-It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon, Brady, and Van Bavel 2022).
-Related to trigger points are dog-whistles—coded expressions that carry two meanings: a seemingly harmless one, and another one with a hidden hateful or provocative connotation that only insiders can decipher (Breitholtz and Cooper 2021).
-In social sciences, the concept of dog whistle has its roots in the study of racial politics in the United States, leading politicians to develop more subtle ways of appealing to racial prejudices (López 2013) and has evolved to encompass a broader range of phenomena, ranging from perpetuated systemic racism (Filimon and and 2024), to more broadly communication strategies of political ambiguity in identity politics(Tolvanen 2024), such as religion (Perry 2023) or immigration (Filimon 2016).
-Trigger points are similar in their equivocality and potential to evoke emotional responses.
-Conversely, using a dog whistle is conscious, but there is typically no intent when relying on trigger points.
-Moreover, while dog whistles are rhetorical devices to hide the real intention from some and create a rally-around-the-flag effect, trigger points are not hidden communication and incite a strong sense of otherness towards those covered by the trigger.
-Research on the automation of identifying affective triggers in online discussions remains limited.
-For instance, Almerekhi et al.
-(2022) developed a framework to identify triggers for toxic behaviour within Reddit discussions.
-Similar work typically uses a restricted set of lexical features, primarily sentiment analysis, and often examines the level of engagement some topics may generate (Aldous, An, and Jansen 2023).
-Others, such as Beel et al.
-(2022), explicitly investigate highly controversial topics like abortion, climate change, and gun control, attempting to predict contentiousness within discussions, often heavily relying on Reddit comment upvote ratios as a proxy for contentiousness.
-Finally, some studies focus on online conversational tensions by predicting conversational derailment, considering the overall progression of a conversation rather than specific terms as causes (Zhang et al.
-2018; Hessel and Lee 2019).
-3 Methodology
-To empirically demonstrate that individual words can trigger online debates, we look into most likely cases in the context of UK politics.
-Based on a curated set of likely candidate words for the political context in the UK (see Section 3.1), we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
-Our goal is to report behavioural mechanisms similar to those observed by Mau, Lux, and Westheuser (2023) when conducting interviews to analyse the perceived polarisation.
-If we could not find these effects in these intuitive cases, we would falsify the theory for online contexts.
-3.1 Term selection
-We selected National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism.
-These terms are associated with high-profile public debates, policy changes, and social movements in the UK.
-Hence, they typically provoke significant public engagement and emotional responses.
-We expect them to cause the intended reactions in British social media users.
-Table 1 indicates mechanisms through which the terms may trigger individuals.
+Studying trigger words helps track the origins of such behaviour, assess whether a social media post can lead to toxic comments (Dahiya et al., 2021), and how language affects polarisation on social media (Karjus and Cuskley, 2024).
+It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
+Trigger words resemble dog whistles—coded expressions with dual meanings, one benign and one provocative to insiders (Breitholtz and Cooper, 2021).
+Originating in US racial politics (López, 2013), the dog whistle concept covers systemic racism (Filimon and and, 2024), political ambiguity in identity discourse (Tolvanen, 2024), religion (Perry, 2023), and immigration (Filimon, 2016).
+Dog whistles are used deliberately to obscure intent.
+Trigger words, however, are not covert but evoke strong affective reactions and heighten perceived otherness.
 Rwanda Brexit NHS Vaccine Feminism
@@
 Behavioural Demands
-Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
+Table 1. Our trigger words and their respective trigger mechanisms (see introduction).
+Research on automating the identification of affective triggers in online discussions remains limited.
+For instance, Almerekhi et al.
+(2022) developed a framework to identify triggers for toxic behaviour within Reddit discussions.
+Similar work typically uses a restricted set of lexical features, primarily sentiment analysis, and often examines the level of engagement some topics may generate (Aldous et al., 2023).
+Others, such as Beel et al.
+(2022), explicitly investigate highly controversial topics, including abortion, climate change, and gun control, attempting to predict contentiousness within discussions.
+They often heavily rely on Reddit comment upvote ratios as a proxy for contentiousness.
+Finally, some studies focus on online conversational tensions by predicting conversational derailment, considering the overall progression of a conversation rather than specific terms as causes (Zhang et al., 2018; Hessel and Lee, 2019).
+3. Methodology
+To empirically demonstrate that individual words can act as triggers in online debates, we look into the most likely cases in the context of UK politics.
+Our goal is to report behavioural mechanisms similar to those observed by Mau et al.
+(2023) offline, i.e., (1) more engagement and (2) animosity.
+If we could not find these effects in these intuitive cases, we would falsify the theory for online contexts.
+3.1. Term selection
+We selected National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism.
+These terms are associated with high-profile public debates, policy changes, and social movements in the UK.
+Hence, they typically provoke significant public engagement and emotional responses, i.e., the intended reactions in British social media users.
+Table 1 suggests mechanisms through which the terms may trigger individuals.
 Rwanda.
-Until late 2021, Rwanda was just the name of any other African country.
+Until late 2021, Rwanda was simply one of many African countries in the eyes of prevailing UK public discourse.
 However, in early 2022, the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership.
-[^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible With this agreement, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist 2023).
+[^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible With this agreement, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
 Brexit.
-The UK’s departure from the European Union has been controversial since its inception.
-The division among the population was already evident during the tightly contested 2016 referendum(Arnorsson and Zoega 2018).
-Additionally, misinformation and disinformation have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022).
-Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from Eastern Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
+The UK’s departure from the European Union has been a contentious issue since its inception.
+The division among the population was already evident during the tightly contested 2016 referendum (Arnorsson and Zoega, 2018).
+Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022).
+Considerations about treating EU migrants differently from the national population (inequality), norm transgressions of “criminal” migrants, or opening the gate to “unlimited” immigration make Brexit a potential trigger word.
 NHS.
-Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the healthcare system (The Guardian 2022; The Independent 2024).
-Public opinion is divided, with some supporting that the NHS receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015).
-This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
+Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the healthcare system (The Guardian, 2022; The Independent, 2024).
+Some support the claim that the NHS receives adequate funding, while others argue that it wastes resources or that its services are in decline (Gershlick et al., 2015).
+This trigger word is related to going against considerations about unequal treatment.
 Vaccine.
-Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years.
-Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021).
-Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al.
-2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
+Recently, public vaccination programs, particularly those against COVID-19, have caused significant debate.
+Post-roll-out, debates have shifted to focus on the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021).
+Moreover, the proliferation of misinformation and conspiracy theories has fueled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media.
 Feminism.
-Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019).
-Some social media influencers have built a large audience by sharing and promoting harmful content against women.
-This has led to the formation of online echo chambers (Farrell et al.
-2019) where such views are amplified and propagated.
-Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope, and unreasonable behavioural demands.
+The anonymity provided by social media often emboldens users to express misogyny (Barker and Jurasz, 2019).
+Some influencers have built a large audience by sharing and promoting harmful content against women.
+This has led to online echo chambers (Farrell et al., 2019) where such views are amplified and propagated.
+We include Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope, and unreasonable behavioural demands.
 Keyword Subreddits
@@
 CoronavirusUK, Coronavirus, wuhan_flu
-Table 2: Subreddits selected for each trigger word.
+Table 2. Subreddits for each trigger word.
 Feminist Rwanda Brexit NHS Vaccine
@@
 Temporal - 49 51 7.2 (2.9, 11.6) - - 50 50 5.9 (4.1, 7.6)
-Table 3: Percentage of comments per thread before (Be) and after (Af) each trigger appearance (Hypothesis 1).
-Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Contextual, comparisons.
+Table 3.
+Percentage of comments per thread before (Be) and after (Af) each trigger appearance (Hypothesis 1).
+Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space, and Contextual comparisons.
 Bold scores indicate statistically significant results.
-Detailed distribution of comments can be seen in Appendix B, see Figure 8.
-3.2 Data collection
-Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al.
-2020).
-We searched for exact matches of specific keywords in comments from 2019 to 2022 and then collected the entire discussion thread for each extracted comment.
+3.2. Data collection
+We collected English-language Reddit comments via the Pushshift API (Baumgartner et al., 2020).
+Searching for exact matches of specific keywords in comments from 2019 to 2022, we collected the entire discussion thread for each extracted comment.
 [^2] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.
-We excluded comments in languages other than English, identified as such using a fastText-based language identifier (Bojanowski et al.
-2017).
-Comments marked as “[deleted]” by the API were also removed.
-Following this, only threads that contained comments before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
-As a final filtering step, we limited our analysis to comments 30 minutes before and 30 minutes after the first mention of the triggering word.
-Previous research has shown that Reddit comments typically receive the majority of interactions within a short window after the comment is created (Thukral et al.
-2018).
-Moreover, early responses are often important in shaping the direction of the conversation (Singer et al.
-2016; Horne, Adali, and Sikdar 2017).
-Considering the fast-paced nature of Reddit, the 30-minute window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data [^3] Table 7 in Appendix B shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered..
-3.3 Research design
-We identify the causal effect of trigger points with the difference-in-difference estimator (DiD), which systematically compares a treatment with a control group (Angrist and Pischke 2009; Card and Krueger 1994).
-It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group.
-In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words.
-There should be a more substantive behavioural change for the treatment group than for the control group.
-Control groups
-We systematically compare treatment and control groups along three different dimensions.
+We excluded comments in languages other than English using a fastText-based language identifier (Bojanowski et al., 2017) and also removed comments that the API marked as “[deleted]”.
+Only threads with comments both before and after the target keywords were retained to enable analysis of pre- and post-trigger differences.
+As a final step, we included only comments posted within 30 minutes before and after the first mention of the trigger word.
+This window captures the most relevant responses, given that Reddit interactions peak quickly and early replies often shape the conversation’s direction (Thukral et al., 2018; Singer et al., 2016; Horne et al., 2017).
+3.3. Research design
+We estimate the causal effect of trigger words using the difference-in-differences (DiD) estimator, which systematically compares a treatment group with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994).
+It calculates the causal effect of an intervention as the difference in the treatment group before and after the intervention minus the difference in the control group before and after the intervention.
+In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms, whereas those in the control group are expected to be much less likely to be triggered by the exact words.
+3.3.1. Control groups
+Discussions on Reddit are organised in different fora (subreddits).
+We systematically compare the treatment and control groups across three dimensions.
 [^3] Identifying the causal effect in three completely unrelated comparisons helps address concerns about eventual endogeneity.
-1. Space
-Discussions on Reddit are structured in different fora, also called subreddits.
-Aiming to investigate the effects a trigger word might have on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 2).
-We compare them with discussions in a control group, i.e., the rest of Reddit with many non-domain specific subreddits.
-2. Contextual
-We also compare trigger words with terms that appear in a similar context, but that does not go against the more profound underlying transgressions responsible for trigger responses in similar contexts.
-To members of the subreddits in Table 2, using an explicit trigger word or its more harmless control term should make a real difference.
-We define related control group words and contrast woman to Feminism/Feminist, Red Cross to NHS, EU/European Union to Brexit, surgery to vaccine, and names of Rwanda’s neighbouring countries [^5] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia.
-to Rwanda.
-Treatment and control groups are from the same subreddits in the same year.
-3. Temporal
-In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time.
-For “Rwanda”, we compare discussions in subreddits likely to spark reactions in 2020 to discussions in 2022.
-For Vaccine we compare threads from 2019 (before the COVID-19 pandemic) to 2022’s (post COVID-19 pandemic).
-Again, we expect that users from the subreddits in Table 2 are triggered by those only in 2022 but not in the respective earlier years.
-Treatment and control groups share the same subreddits but are from different years.
-Statistical analysis
+1.
+Space: To investigate the effects of trigger words on specific subsets of users, we define the treatment group as a set of popular subreddits related to the trigger word (Table 2).
+We compare them with discussions in a control group, i.e., the rest of Reddit, which includes many non-domain-specific subreddits.
+2.
+Contextual: We also compare trigger words to similar terms that lack the deeper transgressive meaning.
+For users in the subreddits listed in Table 2, the difference between a trigger word and a milder control term should make a difference.
+We define control terms by pairing, for example, woman with feminism/feminist, Red Cross with NHS, EU/European Union with Brexit, surgery with vaccine, and Rwanda with its neighbouring countries [^4] Burundi, Tanzania, Congo, Democratic Republic of Congo, Uganda, Kenya, and Zambia..
+Treatment and control data come from the same subreddits and year.
+3.
+Temporal: In the third experiment, we use the fact that two of our terms—Vaccine and Rwanda—only became explicit trigger words at specific points in time.
+For “Rwanda”, we compare discussions in subreddits likely to spark reactions in 2020 with those in 2022.
+For Vaccine, we compare threads from 2019 (pre-COVID-19 pandemic) to 2022 (post-COVID-19 pandemic).
+Again, we expect that users from the subreddits listed in Table 2 were triggered only in 2022, not in earlier years.
+The treatment and control groups share the same subreddits but differ in year.
+3.3.2. Statistical analysis
 We estimate the trigger words’ treatment effect in 60 separate experiments—12 experiments for H1 (Table 3) and 48 experiments for H2 (Table 4).
@@
 We then calculate the share of comments before and after the intervention.
-For example, if there are 8 comments during the 30 minutes before the trigger word and 12 comments during the 30 minutes after the trigger word, we normalize the data of this discussion thread to 40% before intervention and 60% after intervention.
-For experiments with H2, we count the number of comments expressing animosity in a thread and then normalize the data as a share of animosity per thread before and after the trigger word.
-We estimate the causal effect in an OLS-regression framework. More specifically, we specify the DiD with the two-way fixed effects estimator (Angrist and Pischke 2009):
-[…]
-where proportion is the per-thread-normalised target feature and the interaction term is given by
-[…]
-Since the DiD estimator subtracts the difference before and after intervention in the control from the equivalent in the treatment group, no further control variables are necessary—if the parallel trends assumption holds (Roth et al.
-2023).
-It requires that in the absence of treatment, the difference between the outcomes of the treated and control groups would follow the same trajectory over time.
-We investigate the plausibility of the assumption in section 4.3.
-3.4 Features
-We rely on several features to examine our initial hypotheses and study the impact of trigger words in online discussions.
-We use comment counts for the first hypothesis (H1) about user engagement.
-The second hypothesis (H2) requires a measure of animosity.
-To identify relevant features that can operationalize animosity, we build on various language models that follow the RoBERTa architecture (Liu et al.
-2019) and are trained in social media corpora, specifically X (Twitter).
-Even though comments on Reddit tend to be longer than on X, the posts on both platforms share many similarities, e.g., the use of slang, emojis, or unstructured text.
-Hence, we expect the trained models to perform well in either context (Priya et al.
-2019).
-User engagement (Hypothesis 1)
-Our first hypothesis suggests that trigger words lead to higher user engagement.
+For example, if there are 8 comments during the 30 minutes before the trigger word and 12 comments during the 30 minutes after the trigger word, we normalise the data of this discussion thread to 40% before intervention and 60% after intervention.
+For experiments with H2, we count the number of comments expressing animosity in a thread and then normalise the data as a share of animosity per thread before and after the trigger word.
+We identify the causal DiD-estimator with a two-way fixed effects model in an OLS-regression framework (Angrist and Pischke, 2009): $\text{proportion}\sim\text{trigger}+\text{treatment}+\text{interaction}$ where proportion is the per-thread-normalised target feature, and the interaction term is given by $\text{interaction}=\text{trigger}\times\text{treatment}.$ Since the DiD estimator subtracts the difference before and after intervention in the control group from the equivalent in the treatment group, no further control variables are necessary —if the parallel trends assumption holds (Roth et al., 2023).
+In the absence of treatment, the difference in outcomes between the treated and control groups would follow the same trajectory over time.
+[^5] See Section 5 for further explanations and Appendix C for full detailed analyses.
+3.4. Features
+We rely on comment counts to test an increase in user engagement (H1).
+To measure animosity (H2), we identify relevant features using various language models based on the RoBERTa architecture (Liu et al., 2019) that are specialised for social media.
+In Appendix A, we validate the proposed approach through a manual evaluation on a random Reddit sample across a variety of open-source transformer-based language models.
+3.4.1. User engagement (Hypothesis 1)
+Our first hypothesis posits that trigger words increase user engagement.
 We rely on the number of comments to investigate this and expect an uptick in message frequency after the trigger words are mentioned for the first time in a thread.
-Animosity (Hypothesis 2)
-Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalize this with controversial comments, negative sentiment, anger, and different forms of hate speech.
-Controversiality
-We examine whether a comment is controversial using the Reddit controversiality feature.
-Reddit offers a measure that identifies a comment as controversial if it receives similar amounts of upvotes and downvotes.
-This metric is based on the users’ approval, so care must be taken when considering already controversial subreddits.
-We expect trigger words to cause more controversiality.
-Figure 2: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
-Sentiment analysis
-For sentiment analysis, we use twitter-roberta-base-sentiment-latest (Loureiro et al.
-2022) which identifies a comment as negative, neutral, or positive.
-This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal, Farra, and Nakov 2019).
-As we expect trigger words to cause more negative sentiments, we use sentiment analysis to investigate this, similarly to work using it to gain insights into how controversial topics such as vaccination (Melton et al.
-2021), politics (Guimaraes et al.
-2019), and gender biases (Marjanovic, Stańczak, and Augenstein 2022) are discussed online.
-Negative emotions
-We leverage twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al.
-2022) to assign one or more emotions to each tweet.
-This model is trained on the Semeval 2018 Affect in Tweets task (Mohammad et al.
-2018), covering 11 different emotions.
-[^6] Detailed list of emotions can be found in Appendix A.1) For our analysis, we focus on anger specifically, and we expect to observe more anger after a trigger word.
-Figure 3: Distribution of negatively charged comments for the treatment and control (space) groups of Vaccine as a trigger word.
-Hate speech
-Finally, we use the twitter-roberta-base-hate-multi-class model (Antypas and Camacho-Collados 2023) to detect hate speech.
-It is trained on a combination of 13 different hate speech Twitter datasets and can predict hate towards seven target groups.
-[^7] Detailed list of categories can be found in Appendix A.2 ) The inclusion of hate speech detection as a feature is motivated by previous research revealing hateful trends in topics such as politics (Rieger et al.
-2021), women’s rights (Farrell et al.
-2019), and health issues (Walter et al.
-2022).
-Considering the connotations of some of the triggers, we specifically focus on sexism for Feminism/Feminist and racism for Rwanda.
-For NHS, Brexit and Vaccine we consider all hateful comments jointly.
-3.5 Validation of Annotation Model
-The primary aim of our research is to use simple and computationally efficient models that reliably extract signals relevant to our hypotheses.
-To confirm the validity and robustness of our findings, we implemented an additional validation procedure to ensure consistency across the classifiers employed in our analysis.
-We sampled 100 Reddit comments from our data collection, with 20 comments selected from the subreddits associated with each trigger point (Table 2).
-The collected comments were independently annotated by three experienced NLP and CSS researchers for sentiment, emotion, and the presence of hate speech.
-[^8] We include the annotation guidelines in Appendix A.4.
-Each comment was assigned a label based on the majority vote, while disagreements were resolved through discussion.
-When considering the difficulty of the annotation task, the annotations present above-average agreement scores for sentiment and hate speech, with Fleiss’ $\kappa$ scores of 0.52 and 0.56, respectively.
-For the multi-label emotion annotation task, where annotators had to choose from a set of 12 labels,the Krippendorff’s alpha score (Krippendorff 2011) reached 0.24, which is in line with other similar multi-label tasks with a high number of labels (Mohammad et al.
-2018; Antypas et al.
-2022; Ousidhoum et al.
-2019).
-We fine-tune the base version of BERT-cased, XLNet, and RoBERTa for sentiment analysis, emotion recognition, and hate speech detection on the same datasets as the models we used for our experiments, as explained in Section 3.4 (Rosenthal, Farra, and Nakov 2017; Mohammad et al.
-2018; Antypas and Camacho-Collados 2023) to ensure fairness.
-We further compare our results to those of five pre-trained models publicly available on HuggingFace.
-Specifically, two sentiment classifiers trained on Reddit data999 reddit-dBERT: https://huggingface.co/akshataupadhye/finetuning-sentiment-model-reddit-data, reddit-XLNET: https://huggingface.co/minh21/XLNet-Reddit-Sentiment-Analysis , hateBERT (Caselli et al.
-2021) and one emotion classifier [^10] BERT-emotion: https://huggingface.co/ayoubkirouane/BERT-Emotions-Classifier.
-All models are evaluated in the aforementioned manually labelled dataset.
-Table 4 presents the average micro F1-scores achieved by the models across each task.
-We opted for the micro F1-score as it provides a more realistic representation of the results, given the limited sample and heavily imbalanced nature of our dataset.
-[^11] Detailed dataset distribution is available in Appendix A.3 The Twitter-RoBERTa models achieved the highest scores on all tasks.These scores are matched by XLNet-base for sentiment analysis and XLNet for emotion detection, and the fine-tuned models RoBERTa and XLNet for hate speech detection.
-Overall, the tested models performed comparably well, except for two off-the-shelf sentiment classifiers that achieved scores below 50%.
-Sentiment Emotion Hate
-t-RoBERTa 0.76 0.50 0.88
-RoBERTa 0.75 0.47 0.88
-XLNet 0.76 0.45 0.88
-BERT 0.74 0.40 0.85
-reddit-dBERT 0.45 - -
-reddit-XLNet 0.31 - -
-BERT-emotion - 0.45 -
-hateBERT - - 0.80
-Random 0.30 0.20 0.07
-Table 4: Average micro-F1 scores of the models tested for the sentiment analysis, emotion, and hate speech detection tasks. Random indicates a naive baseline that randomly selects labels.
-4 Results
-4.1 User engagement (Hypothesis 1)
-According to our first hypothesis, trigger points cause more comments.
-The DiD estimate reveals significant differences between the frequency of comments in the treatment and control groups for all our trigger words (Table 3).
-In the treatment-only subset, the difference between user engagement before and after the trigger word is apparent for all terms.
-It is the largest with Rwanda where we observe a 10%-points (Table 3) increase.
-To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group.
-In every setting, we observe a significant difference in the number of comments in the threads between the treatment and control subsets, with the frequency of comments in the treatment group increasing by a larger factor according to the DiD estimate.
-The lowest value is a 3.2%-point increase for the contextual comparison in Brexit.
-The largest increase is for NHS in the contextual control group, where we observe a 13.4% point increase.
+3.4.2. Animosity (Hypothesis 2)
+Hypothesis 2 suggests that trigger words cause more animosity in a debate. We operationalise this with controversial comments, negative sentiment, anger, and different forms of hate speech.
+Controversiality. Reddit itself identifies a comment as controversial if it receives similar amounts of upvotes and downvotes. We expect trigger words to cause more controversy.
+Figure 2. Example: Distribution of comments in the treatment and control groups (space) for the trigger word Rwanda.
+Sentiment analysis.
+We use twitter-roberta-base-sentiment-latest (Loureiro et al., 2022) to identify a comment as negative, neutral, or positive.
+This model has been fine-tuned for sentiment analysis using the SemEval 2017 Sentiment Analysis dataset (Rosenthal et al., 2019).
+Our approach builds on existing work that uses sentiment analysis to study online debates on controversial topics such as vaccination (Melton et al., 2021), politics (Guimaraes et al., 2019), and gender bias (Marjanovic et al., 2022).
+Negative emotions.
+Using twitter-roberta-base-emotion-multilabel-latest (Camacho-Collados et al., 2022), we assign one or more emotions to each tweet.
+This model is trained on the Semeval 2018 Affect in Tweets task (Mohammad et al., 2018) and covers 11 different emotions.
+In our analysis, we expect to observe increased anger following a trigger word.
+Hate speech.
+Finally, we use the twitter-roberta-base-hate-multi-class model (Antypas and Camacho-Collados, 2023) to detect hate speech.
+It is trained on a combination of 13 hate-speech Twitter datasets and predicts hate speech directed at seven target groups.
+Hate speech detection is motivated by previous research revealing hateful trends in different topics (Farrell et al., 2019; Rieger et al., 2021; Walter et al., 2022).
+For NHS, Brexit and Vaccine, we consider any hateful comments.
+We focus on sexism for Feminism/Feminist, and on racism for Rwanda.
+4. Results
+4.1. User engagement (Hypothesis 1)
+According to our first hypothesis, trigger words cause more comments.
+The DiD estimate reveals significant differences in comment frequency between the treatment and control groups for all our trigger words (Table 3).
+In the treatment-only subset, the difference in user engagement before and after the trigger word is evident across all terms.
+It is the largest, with Rwanda, where we observe a 10%-point increase (Table 3).
+To identify the causal effect of the trigger word, we compare the difference in the treatment group to the difference in the control group (DiD estimate).
+Across all settings, we observe a significant difference in the number of comments between the treatment and control subsets, with the treatment group showing a larger increase.
+The lowest value is a 3.2 percentage-point increase in the contextual comparison for Brexit.
+The largest increase is for NHS in the contextual control group, with a 13.4% point increase.
 When examining comment frequency across various control groups, the impact of specific trigger words becomes more apparent.
-For instance, Figure 2 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment.
-It shows the overall number of messages across all threads.
+For instance, Figure 2 illustrates the distribution of comments related to the trigger Rwanda in the space-related experiment with the overall number of messages across all threads.
 This trigger word leads to varying levels of engagement across different Reddit communities, with a notable increase in comments in identified communities where Rwanda is more controversial.
@@
 Hate -2.0 (-5.3, 1.3)
-Table 5: Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
-4.2 Animosity (Hypothesis 2)
+Table 4. Difference in Difference (DiD) scores along with the 95% confidence intervals (CI) for each setting tested. Bold scores indicate statistically significant results.
+4.2. Animosity (Hypothesis 2)
 Our second hypothesis suggests that trigger words may increase the frequency of polarised messages and lead to more animosity in the discussion.
-[^12] An example of such a discussion is presented in Appendix E.
-Figure 4: Average controversiality per thread for the treatment and control (space) groups of Rwanda.
+[^6] Examples of discussions are presented in Appendix D.
+Figure 3. Average controversiality per thread for the treatment and control (space) groups of Rwanda.
 Controversiality.
 We analyse the average controversiality of each thread and compare the proportion of controversial comments before and after the introduction of the trigger words.
-Our results (Table 5) indicate that, in general, there is an increase in controversial comments following each trigger word.
-Most terms display a statistically significant increase of DiD estimates when considering controversial messages in the treatment and control groups, with the biggest difference observed on the contextual comparison for Rwanda—a causal effect of 6.6% points.
-However, significant differences in controversiality do not appear across all settings.
-Notably, there seems to be no increase in controversial comments for Feminism, and in the temporal and contextual settings for Vaccine and NHS, respectively.
-Figure 4 shows the percentages of controversial threads related to the trigger word Rwanda for the space experiment, before and after its occurrence.
-The data indicates a noticeable increase in thread controversiality caused by the trigger word, reflected by a DiD estimate of 2.3% points.
-The contextual and temporal control settings follow a similar trend, as documented in Appendix B, Figure 9.
-(a) Rwanda
-(b) Feminism
-Figure 5: Distribution of hateful comments in the treatment and control groups (space) of the trigger points Rwanda and Feminism.
-Sentiment & emotion.
-We further investigate the effect trigger words have on sentiment and emotion.
-Based on our hypothesis, trigger words cause users to express more negatively charged emotions.
-In our analysis, we focus on the presence of negative sentiment and, more narrowly, the presence of anger in emotions.
+Our results (Table 4) indicate that, in general, there are more controversial comments after a trigger word.
+The largest difference is the contextual comparison for Rwanda—a causal effect of 6.6% points.
+However, not all differences are statistically significant.
+Notably, there is no increase in controversial comments for Feminism, and in the temporal and contextual settings for Vaccine and NHS, respectively.
+As an example, Figure 3 shows the percentages of controversial threads related to Rwanda for the space experiment, before and after its occurrence.
+The data indicates a noticeable increase in thread controversiality caused by the trigger word (DiD $2.3\%$ points).
+Sentiment & emotion. Based on our hypothesis, trigger words tend to evoke more negative emotions in users. In our analysis, we focus on negative sentiment, in particular anger.
 Table 4 indicates a clear trend for the negatively charged comments of the treatment and control groups for each trigger word.
 We observe a significant increase in comments conveying a negative sentiment after the trigger word appears in 11 of the 12 settings tested, with the sole exception being the contextual comparison for NHS.
-The largest effect in sentiment, with a DiD estimate of 5.5% points, appears for the Rwanda trigger word in the space and contextual experiments.
-Let us consider the vaccine trigger, Figure 3, as an example.
-Here, we again plot the overall number of messages we observe.
-Both treatment and space control groups exhibit a similar pattern of negative sentiment density as time approaches the trigger word, with an increase in negative sentiment just before and after the trigger word is mentioned.
-However, the treatment group displays a noticeably higher volume of messages around the trigger term, suggesting that triggers amplify the expression of negative sentiment compared to the control condition.
-Similarly to the increase in negative sentiment, trigger words systematically cause more anger.
-We observe a positive causal effect in all conditions, with only 3 non-significant effects.
+The largest effect in sentiment (DiD $5.5\%$ points) appears for the Rwanda trigger word in the space and contextual experiments.
+Trigger words also cause more anger.
+We observe positive causal effects across all conditions, with only 3 non-significant effects.
 Substantively, we measure a causal effect that ranges from 1.4% points more anger (Feminist) to 5.5% points (Vaccine contextual comparison).
-As another example, the overall number of comments expressing anger in the Vaccine treatment and space control groups increase significantly following the appearance of the trigger word, as shown in Figure 12, Appendix B.
-This is further confirmed by the DiD estimates, which indicate a significant difference in anger-related comments in 9 out of 12 cases.
 Hate speech.
@@
 As seen from Table 4, the impact of trigger words regarding the occurrence of hateful or offensive comments is not as prevalent as anger and negative sentiment.
-We observe significant effects in those areas where we can count on a domain-specific hate speech measure, i.e., in the cases of Feminism (space and contextual comparison) and Rwanda (space group).
-We illustrate our findings in a visualization, where we consider all messages irrespective of the thread in which they are posted.
-We observe a distinct pattern, particularly in the context of Rwanda and Feminism, where the volume of hateful content increases after the trigger.
-Figure 5 reveals the prevalence of racist and sexist comments within the treatment groups when compared to their control group in the space experiment.
-4.3 Robustness of Results
-We conduct a series of robustness checks designed to test the sensitivity of our results to assumptions and alternative specifications.
-Testing the Parallel Trends Assumption.
-The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time.
-While there are no formal guarantees for the assumption to hold, there are means to argue in its favour.
-The first argument is a theoretical one.
-In neither of our experimental designs did we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word.
-Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (contextual), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
-We implement event studies for each experiment to further empirically support our argument.
-The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention.
-Plotting the treatment effects at different time points before the treatment, the event study helps visualize whether the pre-treatment trends in the treatment and control groups are indeed parallel.
-If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al.
-2023).
-[…]
-Figure 7 charts the time-wise means over all 60 event studies.
-If the parallel trends assumption holds, the lines before the treatment at time = 0 should all remain on the horizontal line where the treatment effect = 0.
-Indeed, this is largely what we find.
-The mean treatment effect is close to 0 before the trigger word is mentioned for the first time.
-After the intervention, it is clearly above the horizontal 0-line.
-[^13] See Appendix for figures that include individual event studies for all experiments from H1 and H2.
-Placebo Treatment.
-To further validate our Difference-in-Differences (DiD) estimates, we implement a placebo treatment analysis.
-The core logic of this approach is to check whether the estimated treatment effects could be driven by other trends earlier or later than the actual intervention.
-In the placebo analysis, we systematically shift the assumed treatment time in 10-minute intervals, ranging from 30 minutes before to 30 minutes after the actual intervention, i.e., the comment with the first mention of the trigger word in a discussion thread.
-At each placebo point, we re-estimate all DiD models as if the treatment had occurred at that alternative time.
-If our results capture a genuine causal effect, we expect a pronounced treatment effect at the actual intervention time ($t=0$), and near-zero or statistically insignificant estimates at all placebo points before and after.
-Figure 6 visualizes these placebo tests by plotting the precision-weighted average treatment effects and their confidence intervals across all experiments measuring animosity (see Table 5).
-At the actual treatment point $t=0$, which compares message content in the 30 minutes before and after the first appearance of the trigger word ([30,0[[-30,0[[ - 30 , 0 [ vs.
-$[0,30]$), we observe a clear and substantial effect (1.99).
-We then shift the treatment window rightward in 10-minute steps—first comparing [20,10[[-20,10[[ - 20 , 10 [ to $[10,40]$, and so on—until the actual intervention fully exits the analysis window at $t=30$.
-This generates a steady decline in estimated effects, which converge to zero, as expected if no real treatment occurred at these placebo times.
-Similarly, shifting the window to earlier times (e.g., comparing [60,30[[-60,-30[[ - 60 , - 30 [ to $[-30,0]$) reveals a decreasing trend in effect size, though some residual difference remains.
-In conclusion, given the clear peak of the effect at the timing $t=0$, we are confident in the robustness of our approach to defining the moment of the actual treatment moment.
-Figure 6: Placebo treatments every 10 minutes, ranging from 30 minutes before to 30 minutes after the actual treatment point.
-Precision weighted average treatment effects and 95% confidence intervals across all experiments.
-Figure 7: Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
-5 Conclusion
-This paper establishes trigger words as a concept in online debates.
-Since our work represents the first systematic computational study of this concept, our research design and case selection are geared to clearly identify such behaviour in the most likely cases.
-To do so, we performed a large-scale analysis about UK politics on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS, and Vaccine).
-The results show that trigger words cause more user engagement, and this activity is marked by higher levels of animosity.
-This confirms the first hypothesis (H1), which suggests more user interaction after the appearance of a trigger word.
-Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects.
+We observe significant effects in areas where we can rely on a domain-specific hate speech measure, i.e., in Feminism (space and contextual comparison) and Rwanda (space group).
+5. Robustness of Results
+We assess the robustness of our Difference-in-Differences (DiD) estimates through parallel trends tests and placebo treatments.
+The parallel trends assumption requires that, absent treatment, the treatment and control groups would follow similar trends over time.
+To assess this, we implement event studies that trace changes in the outcome variable across six 10-minute intervals before and after the first mention of a trigger word.
+As shown in Appendix C, Figure 6, pre-treatment trends remain flat and centred around zero, supporting the validity of the DiD design.
+To further validate our identification strategy, we conduct placebo analyses by shifting the assumed treatment point in 10-minute increments before and after the actual intervention.
+For each placebo word, we re-estimate the DiD models.
+If our findings were artefacts of general conversational trends, we would expect similar effects with these alternative words.
+Instead, Figure 5 in Appendix C shows that only the actual treatment time yields a pronounced effect, while estimates at placebo times converge toward zero.
+Together, our checks strengthen confidence in our causal claims.
+6. Conclusion
+Our work represents the first systematic computational study of trigger words in online debates.
+The research design and case selection are geared toward identifying such behaviour in the most likely cases, namely a large-scale analysis of UK politics on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS, and Vaccine).
+The results show that trigger words lead to more user engagement and higher levels of animosity.
+This confirms the first hypothesis (H1), which predicts increased user interaction upon encountering a trigger word.
+Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity in ensuing discussions, we examine four features that may indicate such effects.
 Trigger words cause more controversiality, more negative sentiment, and more anger.
-We also investigate the causal effect of trigger words on hate speech.
-Where we can, we rely on a model capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—.
-Our findings similarly confirm that trigger words play a causal role in this aspect as well.
-From this work, we can certify that trigger points exist in online communication, and we have provided a framework to measure the extent of them in terms of user engagement and animosity.
+Our findings also confirm that trigger words cause more hate speech.
+From this work, we can certify that trigger words exist in online communication.
 Future work will analyse this phenomenon more comprehensively.
-We aim to broaden the scope, study how prevalent trigger points are in different countries and languages, and seek to better understand further behavioural consequences from trigger points.
-It is our intention to evaluate the extent to which trigger points are responsible for hateful online communication.
-We also strive to automate the trigger point detection process to detect the respective words in online conversations when we do not know them ex-ante.
-6 Limitations
-We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset ($>$100 million entries).
+We aim to broaden the scope, study the prevalence of trigger words across countries and languages, and better understand their behavioural consequences.
+We also strive to automate trigger-word detection to identify relevant words in online conversations when we are not familiar with them a priori.
+7. Limitations
+We refrain from using large language models (LLMs) for feature extraction primarily because of the size of our dataset (>>100 million entries).
 Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments.
-Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al.
-2023).
-Our study does not attempt to map how pervasive trigger points are in online debates.
-The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community.
-Similarly, we acknowledge the limitations and potential biases of our work, as we only investigate a small set of potential trigger words in the context of the UK and limit our data to comments written in English in a limited selection of subreddits.
-The choice of trigger words is informed by a contextual analysis of the UK politics, and the Reddit threads have been selected due to the often polarised conversations they host around such topics.
-However, this is the first study that analyses trigger points’ concept validity in online conversations.
-Based on our empirical evidence, we can be confident that we have found relevant evidence that warrants further investigation.
-7 Ethics Statement
+Moreover, the existing literature indicates that LLMs’ zero-shot capabilities do not consistently outperform smaller models on specific social media tasks such as ours (Antypas et al., 2023).
+Our study does not aim to map the prevalence of trigger words across online debates.
+The selected subreddits may not fully represent the broader range of discussions these trigger words could elicit across the Reddit community.
+Similarly, we acknowledge the limitations and potential biases of our work, as we investigate only a small set of trigger words in the UK and limit our data to English-language comments from a limited selection of subreddits.
+However, as the first study to analyse the validity of trigger words in online discussions, we are confident that our findings warrant further investigation.
+8. Ethics Statement
 All data used in our research is sourced from publicly available information or collected using the official Reddit API.
-It is provided in an aggregated way and without reporting sensitive information from individual users.
+We present aggregated data without reporting sensitive information from individual users.
 The automatic methods used can be abused by those with the power to repress people with opposing opinions.
-Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use.
-Researchers interested in replicating our experiments can contact us directly.
+Hence, we provide a comprehensive study but do not share our data publicly to prevent potential misuse or commercial exploitation.
+Researchers interested in replicating our experiments are encouraged to contact us directly.
 Appendix A Model Selection
-A.1 Emotion Categories
-The twitter-roberta-base-emotion-multilabel model classifies each entry in one or more of the following classes: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust.
-A.2 Hate Speech Categories.
-The twitter-roberta-base-hate-multiclass model classifies each entry in one of the following classes: not_hate, sexism, racism, religion, other, sexual_orientation, disability
-A.3 Validation Setting
-Table 6 displays the distribution of comments in regards to emotion, sentiment and hate speech for the manually annotated dataset collected.
+A.1. Validation of Annotation Model
+We use computationally efficient models that can be deployed locally to extract signals relevant to our hypotheses, and then validate the annotation procedure.
+We sampled 100 Reddit comments from our dataset, 20 for each trigger word, and then hand-annotated sentiment, emotion, and the presence of hate speech with three experienced NLP and CSS researchers.
+Each comment was assigned a label by majority vote, and disagreements were resolved through discussion (see Table 5).
+Given the difficulty of the annotation task, the annotations above show above-average agreement scores for sentiment and hate speech, with Fleiss’ $\kappa$ scores of 0.52 and 0.56, respectively.
+For the multi-label emotion annotation task, where annotators had to choose from a set of 12 labels, the Krippendorff’s alpha score (Krippendorff, 2011) reached 0.24, which is in line with other similar multi-label tasks with a high number of labels (Mohammad et al., 2018; Antypas et al., 2022; Ousidhoum et al., 2019).
 Emotion Hate Speech Sentiment
@@
 joy 2
-Table 6: Distribution of comments regarding emotion, sentiment, and hate speech categories on the manually annotated dataset.
-A.4 Annotation Guidelines
-Below we present the guidelines presented to annotators for the Sentiment Analysis, Emtion classification, and Hate Speech detection tasks.
-Sentiment Analysis
-We will show you Reddit comments that can be positive, negative, or neutral based on their sentiment.
-Sentiment Categories
-  • Positive comments should be labelled: Positive
-  • Neutral comments should be labelled: Neutral
-  • Negative comments should be labelled: Negative
-Examples
-  • “‘Just finished my morning workout and feeling fantastic! #HealthyLiving #FitnessGoals“‘ is a positive sentence.
-  • “‘Stuck in traffic again! Late for work and already stressed out. #TrafficJam“‘ #MondayBlues is a negative sentence.
-  • Factual sentences such as “‘The party was in front of the house“‘ are neutral sentences.
-Notes: We have also added a comment column where you can share your thoughts, comments, or doubts.
-If you are unsure whether you should annotate the tweet as positive, negative or neutral, you may choose the label that suits it best and a comment about that in the comment column.
-The annotation of an ironic comments depends on the context. While we acknowledge the lack of context, if you think the tweet is mocking someone (or a group of people), you should label it negative.
-Emotion Classification
-You will be presented with Reddit comments and asked to select the option that best describes the text.
-You may choose as many emotions as you believe are present — from 0 to 11 — based on the content and tone of the comment.
-Emotion Categories:
-  • Anger – includes annoyance, rage
-  • Anticipation – includes interest, vigilance
-  • Disgust – includes disinterest, dislike, loathing
-  • Fear – includes apprehension, anxiety, terror
-  • Joy – includes serenity, ecstasy
-  • Love – includes affection
-  • Optimism – includes hopefulness, confidence
-  • Pessimism – includes cynicism, no confidence
-  • Sadness – includes pensiveness, grief
-  • Surprise – includes distraction, amazement
-  • Trust – includes acceptance, liking, admiration
-Notes: If no emotion is clearly present, you may leave all options unchecked.
-Choose all emotions that are clearly expressed or implied in the comment.
-If you’re unsure, use the comment column to share your reasoning or highlight ambiguity.
-Comments may express multiple emotions simultaneously, for example, both sadness and anger.
-Hate Speech Detection
-You will be presented with Reddit comments and asked to identify whether they contain hate speech, and if so, which target group is being attacked or discriminated against.
-Please choose only one of the following categories that best represents the nature of the content:
-Hate Speech Categories:
-  • Sexism: Content that targets or demeans people based on their gender (e.g., women, men, non-binary individuals) or reinforces harmful gender stereotypes.
-Example: “Women should stay in the kitchen where they belong.”
-  • Racism: Content that expresses hostility, prejudice, or dehumanization based on race, ethnicity, or national origin. Example: “All [ethnic group] are criminals and should be deported.”
-  • Disability: Content that mocks, belittles, or degrades people with physical, mental, or developmental disabilities. Example: “You’re so stupid, you must be on the spectrum.”
-  • Sexual Orientation: Hate directed toward someone because of their actual or perceived sexual orientation (e.g., gay, lesbian, bisexual, etc.).
-Example: “Being gay is a disease that should be cured.”
-  • Religion: Disparagement, dehumanization, or threats against individuals or groups based on religious beliefs or identities. Example: “All Muslims are terrorists.”
-  • Other: Use this label for hate speech that targets a group not covered above, such as socioeconomic status, age, appearance, political affiliation, etc.
-Example: “Old people are useless and should just die already.”
-  • Not Hate: Use this label if the comments does not contain hate speech. Comments that are neutral or non-discriminatory.
-Notes: If you are unsure, select the label you think fits best and add a comment explaining your reasoning.
-A single post may mention multiple groups in such a case, select the one that is most explicit or harmful, and explain in the comment column.
+Table 5. Distribution of comments regarding emotion, sentiment, and hate speech categories on the manually annotated dataset.
+We fine-tune the base version of BERT-cased, XLNet, and RoBERTa for sentiment analysis, emotion recognition, and hate speech detection on the same datasets as the models we used for our experiments (Rosenthal et al., 2017; Mohammad et al., 2018; Antypas and Camacho-Collados, 2023).
+We further compare our results against five publicly available pre-trained models on Hugging Face.
+Specifically, two sentiment classifiers trained on Reddit data777reddit-dBERT: https://huggingface.co/akshataupadhye/finetuning-sentiment-model-reddit-data, reddit-XLNET: https://huggingface.co/minh21/XLNet-Reddit-Sentiment-Analysis , hateBERT (Caselli et al., 2021) and one emotion classifier [^8] BERT-emotion: https://huggingface.co/ayoubkirouane/BERT-Emotions-Classifier.
+All models are evaluated in the aforementioned manually labelled dataset.
+Table 6 presents the average micro F1-scores achieved by the models across each task.
+[^9] Micro F1-score provide a more realistic representation of the results, given the limited sample and heavily imbalanced data.
+The Twitter-RoBERTa models achieved the highest scores on all tasks.
+These scores are matched by XLNet-base for sentiment analysis and XLNet for emotion detection, and the fine-tuned models RoBERTa and XLNet for hate speech detection.
+Overall, the tested models performed comparably well, except for two off-the-shelf sentiment classifiers that achieved scores below 50%.
+Sentiment Emotion Hate
+t-RoBERTa 0.76 0.50 0.88
+RoBERTa 0.75 0.47 0.88
+XLNet 0.76 0.45 0.88
+BERT 0.74 0.40 0.85
+reddit-dBERT 0.45 - -
+reddit-XLNet 0.31 - -
+BERT-emotion - 0.45 -
+hateBERT - - 0.80
+Random 0.30 0.20 0.07
+Table 6. Average micro-F1 scores of the models tested for the sentiment analysis, emotion, and hate speech detection tasks. Random indicates a naive baseline that randomly selects labels.
+A.2. Annotated Categories
+  • Emotion Categories: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust
+  • Hate Speech Categories: not_hate, sexism, racism, religion, other, sexual_orientation, disability
 Appendix B Extended Results
-Table 7 presents the total number of comments and threads extracted from Reddit for each keyword, along with the corresponding time periods during which the data was collected.
 Table 7 presents the number of comments included in each experiment for both the treatment and control groups.
-Figure 8 displays the number of comments posted before and after the introduction of each trigger point on the data collected.
-Figure 9 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
-Figure 10 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
-Figure 11 displays the distribution of negative comments, based on their sentiment, posted before and after the introduction of the Brexit, Feminism, Rwanda, and NHS trigger points in the Space control setting.
-Figure 12 displays the distribution of comments conveying anger and those that do not before and after the introduction of the Vaccine trigger point in the Space control setting.
-Threads Comments Year
-Rwanda 0.05 1.87 2020, 2022
-Brexit 0.37 11.57 2020 - 2022
-NHS 0.34 6.20 2020 - 2022
-Vaccine 1.42 28.11 2019, 2022
-Feminism 0.94 17.79 2020 - 2022
-EU 0.2 12.4 2020 - 2022
-Red Cross 0.02 0.2 2020 - 2022
-SubSaharan 0.08 1.5 2020, 2022
-Surgery 0.5 3.9 2019, 2022
-Woman 2.1 19.3 2020 - 2022
-Table 7: Numbers, in millions, of threads, comments, and time period for each trigger word.
+Figure 4 displays the distribution of comments posted before and after the introduction of the Rwanda trigger word in the Temporal and Contextual control settings.
 Trigger Experiment Treatment Control
@@
 contextual 1,276
-Table 8: Number of comments present for each trigger in each experiment for both the treatment and control subsets.
-(a) Brexit
-(b) Feminism
-(c) NHS
-(d) Rwanda
-(e) Vaccine
-Figure 8: Absolute count of comments before and after each trigger point.
-(a) Temporal Setting
-(b) Contextual Setting
-Figure 9: Average controversiality per thread in the Temporal and Contextual settings and their respective treatment groups, for the trigger point Rwanda.
+Table 7. Number of comments present for each trigger in each experiment for both the treatment and control subsets.
 (a) Contextual Comparison
 (b) Temporal Comparison
-Figure 10: Rwanda Treatment and Control distributions of comments in the Contextual and Temporal settings.
-(a) Brexit
-(b) Feminism
-(c) Rwanda
-(d) NHS
-Figure 11: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
-Figure 12: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
-Appendix C Extended Test of Parallel Trends Assumption
-Figures 13 a & b display the results of the Parallel Trends Assumption testing for all experiments as described in H1 and H2.
-(a) Experiments from H1.
-(b) Experiments from H2.
-Figure 13: Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
-Appendix D Computational Resources
-An NVIDIA GeForce RTX 4090 GPU was utilised for the experiments conducted. 170 hours spent for the inference process of sentiment analysis, emotion detection, and hate speech detection models.
-Appendix E Selected Reddit Comments
-Below we present a Reddit discussion and how it evolves when the trigger ”feminism” appears. The comment containing the trigger is bolded.
+Figure 4. Rwanda Treatment and Control distributions of comments in the Contextual and Temporal settings.
+Appendix C Robustness of Results
+We conduct a series of robustness checks designed to test the sensitivity of our results to assumptions and alternative specifications.
+Testing the Parallel Trends Assumption.
+The DiD estimator relies on the assumption of parallel trends.
+While there are no formal guarantees that the assumption holds, there are grounds to argue in its favour.
+The first argument is theoretical.
+In neither of our experimental designs did we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word.
+Whether comparing treatment and control groups across different Reddit subreddits (spaces), different words (contextual), or points in time (temporal), without the trigger word, the relative difference between treatment and control groups would remain the same.
+To further empirically support our argument, we conduct event studies.
+For each experiment, we model the outcome variable over time, enabling us to assess the dynamics of treatment impact before and after the intervention.
+Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel (Roth et al., 2023).
+For each of our 60 overall experiments, we divide the 30 minutes before and after the first mention of a trigger word into six 10-minute intervals, rather than just two.
+This allows us to track pre-treatment trends at higher granularity.
+We calculate each discussion thread’s share of (hostile) messages in the six intervals.
+Averaging across all data points per period, we subtract the control group from the treatment group and compute the difference for each observation relative to the first observation period.
+Figure 6 charts the time-wise means over all 60 event studies.
+If the parallel trends assumption holds, the lines before the treatment at time =0=0 should all remain on the horizontal line where the treatment effect =0=0.
+Indeed, this is essentially what we find.
+The mean treatment effect is close to 0 before the trigger word is first mentioned.
+After the intervention, it is clearly above the horizontal 0-line.
+Figure 5. Placebo treatments every 10 minutes, ranging from 30 minutes before to 30 minutes after the actual treatment point.
+Placebo Treatment.
+We also implement a placebo treatment analysis.
+The core logic of this approach is to check whether the estimated treatment effects could be driven by other trends earlier or later than the actual intervention.
+Shifting the assumed treatment time in 10-minute intervals, ranging from 30 minutes before to 30 minutes after the actual intervention, we re-estimate all DiD models at each placebo point.
+Figure 5 visualises these placebo tests by plotting the precision-weighted average treatment effects and their confidence intervals across all experiments measuring animosity (see Table 4).
+At the actual treatment point $t=0$, which compares message content in the 30 minutes before and after the first appearance of the trigger word ([30,0[[-30,0[ vs.
+$[0,30]$), we observe a clear and substantial effect (1.99).
+We then shift the treatment window rightward in 10-minute steps—first comparing [20,10[[-20,10[ to $[10,40]$, and so on—until the actual intervention fully exits the analysis window at $t=30$.
+This generates a steady decline in estimated effects, which converge to zero, as expected if no real treatment occurred at these placebo times.
+Similarly, shifting the window to earlier times (e.g., comparing [60,30[[-60,-30[ to [30,0][-30,0]) reveals a decreasing trend in effect size, though some residual difference remains.
+In conclusion, given the clear peak in the effect at $t=0$, we are confident in the robustness of our approach to defining the actual treatment moment.
+Figure 6. Testing the Parallel Trends Assumption with event studies.
+Appendix D Selected Reddit Comments
+Below, we present a Reddit discussion and how it evolves when the trigger “feminism” is used. The comment containing the trigger is bolded.
 WARNING: the following comments contain sensitive and potentially offensive language. Reader discretion is advised.